### v2

# Understanding the Effectiveness of Large Language Models in Detecting Security Vulnerabilities

Abstract.
Security vulnerabilities in modern software are prevalent and harmful. While automated vulnerability detection tools have made promising progress, their scalability and applicability remain challenging. Recently, Large Language Models (LLMs), such as GPT-4 and CodeLlama, have demonstrated remarkable performance on code-related tasks. However, it is unknown whether such LLMs can do complex reasoning over code. In this work, we explore whether pre-trained LLMs can detect security vulnerabilities and address the limitations of existing tools. We evaluate the effectiveness of pre-trained LLMs, in terms of performance, explainability, and robustness, on a set of five diverse security benchmarks spanning two languages, Java and C/C++, and covering both synthetic and real-world projects.
Overall, all LLMs show modest effectiveness in end-to-end reasoning about vulnerabilities, obtaining an average of 60% accuracy across all datasets. However, we observe that LLMs show promising abilities at performing parts of the analysis correctly, such as identifying vulnerability-related specifications (e.g., sources and sinks) and leveraging natural language information to understand code behavior (e.g., to check if code is sanitized). Further, LLMs are relatively much better at detecting simpler vulnerabilities that typically only need local reasoning (e.g., Integer Overflows and NULL pointer dereference). We find that advanced prompting strategies that involve step-by-step analysis significantly improve performance of LLMs on real-world datasets (improving F1 score by up to 0.25 on average). Finally, we share our insights and recommendations for future work on leveraging LLMs for vulnerability detection.
1. Introduction
Security vulnerabilities afflict software despite decades of advances in programming languages, program analysis tools, and software engineering practices. Even well-tested and critical software such as OpenSSL, a widely used library for applications that provide secure communications, contains trivial buffer overflow vulnerabilities, e.g., (CVE-2022-3602, 2022) and (CVE-2022-3786, 2022). A recent study by Microsoft showed that more than 70% of the vulnerabilities are still caused by well-understood memory safety issues (Miller, 2019). The size and complexity of modern software systems are growing quickly, encompassing numerous programs, libraries, and modules that interact with each other. Hence, we need major technical advances to effectively detect security vulnerabilities in such complex software.
Traditional techniques for automated vulnerability detection, such as fuzzers (Manes et al., 2018), and static analyzers such as CodeQL (Avgustinov et al., 2016) and Semgrep (Semgrep, 2023) have made promising strides. For example, in the last two years, researchers found over 300 security vulnerabilities through custom CodeQL queries (Semmle, 2023; Leong, 2022). However, these techniques face challenges in scalability and applicability. Fuzzing does not scale to large applications, and fuzzing parts of applications requires manually crafting fuzz drivers. Moreover, it is hard to use on large critical programs with complex inputs, such as network servers, GUI-based programs, embedded firmware, boot loaders, and system services. On the other hand, static analysis relies heavily on manual API specifications, and skillfully crafted heuristics to balance precision and scalability. In light of these challenges, GitHub pays a bounty of over 7K USD for each CodeQL query that can find new critical security bugs (GitHub, 2023).
Large Language Models (LLMs), including pre-trained models such as GPT-4 and CodeLlama, have made remarkable advances in code-related tasks in a relatively short period. Such tasks include code completion (Chen et al., 2021), automated program repair (Xia et al., 2023; Joshi et al., 2023; Xia and Zhang, 2022), test generation (Lemieux et al., 2023; Deng et al., 2023), code evolution (Zhang et al., 2023), and fault localization (Yang et al., 2024). These results clearly show the promise of LLMs, opening up a new direction for exploring advanced techniques. Hence, an intriguing question is whether the state-of-the-art pre-trained LLMs can also be used for detecting security vulnerabilities in code.
To develop LLM-based solutions, an important first step is to systematically evaluate the ability of LLMs in detecting known vulnerabilities. This is especially important in light of the rapidly evolving landscape of LLMs in three aspects: scale, diversity, and applicability. First, scaling these models to ever larger numbers of parameters has led to significant improvements over previous generations in their capabilities—a phenomenon termed as emergent behavior (Wei et al., 2022a). For instance, GPT-4, which is presumably orders of magnitude larger than its 175-billion predecessor GPT-3.5, significantly outperforms GPT-3.5 on a wide range of code-understanding tasks (Bubeck et al., 2023). Second, the diversity of LLMs has grown rapidly and now includes not only proprietary general-purpose ones such as GPT-4 but also open-sourced LLMs such as CodeLlama (Rozière et al., 2023) and StarCoder (Li et al., 2023) that are specialized for code. Finally, the reasoning capabilities of LLMs (and hence their applicability) may vary significantly across different programming languages. All these factors open up a large exploration space for applying LLMs to the challenging task of vulnerability detection.
Our Work. We conduct the first comprehensive study of using LLMs for detecting security vulnerabilities. We study five state-of-the-art LLMs, including proprietary models such as GPT-3.5 and GPT-4, and open-source models like CodeLlama. We evaluate these models on five popular security vulnerability datasets.
We design a set of three prompting strategies for LLMs to elicit increasingly sophisticated forms of reasoning and explanations. Our simplest prompting strategies include the Basic prompt, which simply asks an LLM to check for any vulnerabilities in the given code and the CWE specific prompt, which asks the LLM to check for a specific class of vulnerabilities or CWEs (such as Buffer Overflows).
A significant limitation of static vulnerability detection tools is the requirement of building the target project to enable them to find bugs. Further, they also require concrete specifications of APIs (e.g., sources, sanitizers, and sinks). In contrast, LLMs have an internal model of APIs already seen during the pre-training phase and hence they do not require compiled or complete codebases to run. Inspired by this insight, we additionally design a new prompting strategy, called Dataflow analysis-based prompt, that simulates a source-sink-sanitizer based dataflow analysis on the target code snippet before predicting if it is vulnerable. The Dataflow analysis-based prompt, similar to a classical dataflow-based static analyzer asks the LLM to first infer the sources, sinks, and sanitizers in the code snippet and check for any unsanitized data flows between the source and sink. This style of prompting is similar to chain-of-thought reasoning and simulates step-by-step reasoning by LLMs.
A key challenge in evaluating LLMs is data leakage: LLMs may perform well on a dataset because such samples were already present in their pre-training data. Hence, we implement three semantics-preserving adversarial attacks for code and evaluate whether they significantly impact the performance of LLMs.
Research Questions. We study the following research questions:
RQ1: How do different pre-trained LLMs perform in detecting security vulnerabilities across different languages and datasets? (Section §3.1)
RQ2: How do different prompting strategies affect the performance of LLMs? (Section 3.2)
RQ3: How does the performance of LLMs vary across different vulnerability classes? (Section 3.3)
RQ4: Can adversarial attacks impact the performance of LLMs in detecting vulnerabilities? (Section §3.4)
RQ5: How do LLMs compare to state-of-the-art static analysis tools? (Section 3.5)
RQ6: How do LLMs compare to state-of-the-art deep-learning-based tools? (Section 3.6)
Results and Findings. We choose five state-of-the-art LLMs: GPT-4, GPT-3.5, CodeLlama-34B, CodeLlama-13B, and CodeLlama-7B and evaluate their effectiveness in detecting security vulnerabilities across five vulnerability datasets: OWASP (OWASP Benchmark Suite, 2023), Juliet Java (Juliet Java, 2023), Juliet C/C++ (Juliet C/C++, 2023), CVEFixes Java (Bhandari et al., 2021), and CVEFixes C/C++ (Bhandari et al., 2021). These datasets cover 5000 vulnerable and non-vulnerable code samples, across 25 CWEs. Our findings are summarized as follows:
LLMs across all sizes have modest vulnerability detection ability, with an average accuracy of only about 60% across all datasets.
Using prompts that focus on detecting specific CWEs improves the performance of LLMs. The dataflow analysis-based prompt further improves results for larger LLMs by allowing step-by-step reasoning. The improvement is significant for real-world datasets (an increase of up to 0.25 F1 score). However, we observe that LLMs often infer the relevant source, sink, and sanitizers correctly but fail in end-to-end logical reasoning.
LLMs are relatively better at detecting vulnerabilities that require local reasoning without complex code structures (such as Integer Overflow and NULL Pointer Dereference). However, they struggle to detect more complex vulnerabilities in real-world programs.
Larger LLMs like GPT-4 show mild degradation in performance when subject to adversarial attacks on code (average drop 8.6%). In contrast, smaller LLMs, like CodeLlama-7B, show a significant drop in performance (average drop 39%) in the presence of perturbations.
LLMs lag behind state-of-the-art static analysis tools like CodeQL in average accuracy (by 15%) and precision (by 26%) across datasets. However, LLMs like GPT-4 show some promising partial abilities, e.g., by identifying source/sink specifications or doing context-based reasoning using natural language information, which can potentially be useful for static analysis.
We obtain mixed results comparing LLMs to prior deep-learning-based tools: while the deep-learning-based tools only slightly outperform LLMs on synthetic datasets (by 0.07 in F1 score), all approaches struggle equally on real-world datasets often producing close to 50% accuracy only (same as a random baseline).
Contributions. To summarize, we make the following contributions in this paper:
Empirical Study: We conduct the first large comprehensive study on how state-of-the-art LLMs perform in detecting security vulnerabilities across five datasets and two programming languages (C/C++ and Java).
Prompting Strategies: We design three prompting strategies for LLMs, inspired by the recent advances in natural language processing and traditional program analysis techniques, that elicit different reasoning styles from LLMs and also provide human-readable explanations for their predictions.
Robustness of LLMs: We study how the performance of LLMs is impacted by semantics-preserving adversarial attacks on code.
Comparison with other vulnerability detection tools: We contrast the performance of LLMs against popular static analysis and deep-learning-based vulnerability detection tools.
Insights: We perform a rigorous manual analysis of LLMs’ predictions and highlight vulnerability patterns that impact the performance of these models.
2. Approach
2.1. Datasets
For our study, we select five diverse vulnerability datasets from two languages: C++ and Java. For each language, we also select both synthetic and real-world benchmarks. Table 1 presents the details of each dataset, such as the dataset size, programming language, number of vulnerable and non-vulnerable samples, and the number of unique CWEs. We describe each dataset next.
| Dataset | Language | Size | Vul/Non-Vul | CWEs | |
| OWASP(OWASP Benchmark Suite, 2023) | Java | 2740 | 1415/1325 | 11 | |
| SARD Juliet (C/C++) (Juliet C/C++, 2023) | C/C++ | 81,280 | 40,640/40,640 | 118 | |
| SARD Juliet (Java) (Juliet Java, 2023) | Java | 35,940 | 17,970/17,970 | 112 | |
| CVEFixes (Bhandari et al., 2021) | C/C++ | 19,576 | 8223/11,347 | 131 | |
| CVEFixes (Bhandari et al., 2021) | Java | 3926 | 1461/2465 | 68 |
2.1.1. OWASP (Synthetic)
The Open Web Application Security Project (OWASP) benchmark (OWASP Benchmark Suite, 2023) is a Java test suite designed to evaluate the effectiveness of vulnerability detection tools. Each test represents a synthetically designed code snippet containing a security vulnerability. OWASP contains 2740 test cases representing 11 unique classes of security vulnerabilities (also known as Common Weakness Enumeration or CWE).
2.1.2. Juliet (Synthetic)
Juliet (Boland and Black, 2012) is a widely-used vulnerability dataset developed by NIST. Juliet comprises thousands of synthetically generated test cases representing various known vulnerability patterns. It contains 81,280 C/C++ programs covering 118 unique CWEs, and 35,940 Java programs covering 112 unique CWEs. For our paper, we use the latest version, Juliet 1.3 (Black and Black, 2018; Juliet C/C++, 2023; Juliet Java, 2023).
2.1.3. CVEFixes (Real-World)
Bhandari et al. (Bhandari et al., 2021) curated a dataset, known as CVEFixes, from 5365 Common Vulnerabilities and Exposures (CVE) records from the National Vulnerability Database (NVD). From each CVE, they automatically extracted the vulnerable and patched versions of each method in open-source projects, along with extensive meta-data such as the corresponding CWEs, project information, and commit data. CVEFixes consists of methods extracted from 5495 vulnerability-fixing commits. These methods span multiple programming languages such as C/C++, Java, Python, and JavaScript. For our work, we extracted all C/C++ and Java methods from CVEFixes. We collected 19,576 C/C++ and 3926 Java methods (both vulnerable and non-vulnerable), covering 131 and 68 different CWEs, respectively.
While many real-world datasets have been proposed in the literature, we needed a dataset for our study that 1) contains vulnerability metadata such as CVE and CWE IDs, 2) is two-sided, i.e., contains both vulnerable and non-vulnerable code samples, and 3) covers multiple languages such as Java and C/C++. Table 2 shows a comparison of existing real-world vulnerability datasets. We selected CVEFixes because it is the only dataset that fits our criteria.
| Dataset | Languages | CVE Metadata | Two-Sided | Multi-Lang |
| BigVul (Fan et al., 2020) | C/C++ | ✓ | ✗ | ✗ |
| Reveal (Chakraborty et al., 2021) | C/C++ | ✗ | ✗ | ✗ |
| DiverseVul (Chen et al., 2023) | C/C++ | ✗ | ✓ | ✗ |
| DeepVD (Wang et al., 2023) | C/C++ | ✗ | ✗ | ✗ |
| CVEFixes (Bhandari et al., 2021) | C/C++, Java, … | ✓ | ✓ | ✓ |
2.2. Metrics
To evaluate the effectiveness of each tool, we use the standard metrics used for classification problems. In this work, a true positive represents a case when a tool detects a true vulnerability. In contrast, a false positive is when the tool detects a vulnerability that is not exploitable. True and false negatives are defined analogously. We describe each metric in the context of vulnerability detection.
Accuracy: Accuracy measures how often the tool makes a correct prediction, i.e., whether a code snippet is vulnerable or not. It is computed as: .
Precision: Precision represents what proportion of cases that a tool detects as a vulnerability is a correct detection. It is computed as: .
Recall: Recall represents what proportion of vulnerabilities the tool can detect. It is computed as: .
F1 score: The F1 score is a harmonic mean of precision and recall. It is computed as:
2.3. Large Language Models
We choose the most popular state-of-the-art pre-trained Large Language Models (LLMs) for our evaluation. We choose OpenAI models: GPT-4 (gpt-4) and GPT-3.5 (gpt-3.5-turbo). GPT-3.5 allows up to 4096 input tokens while GPT-4 (which presumably is much larger) allows up to 8192 tokens in the input prompt. Since these models are closed-source, we also evaluate CodeLlama models (Rozière et al., 2023), which were recently open-sourced by Meta. We select three versions of these models: CodeLlama-7B (CL-7B), CodeLlama-13B (CL-13B), and CodeLlama-34B (CL-34B), containing 7 billion, 13 billion, and 34 billion parameters, respectively. We use the Hugging Face APIs (Hugging Face, 2023) to access CodeLlama models. We use the “Instruct” version of CodeLlama models—these models are fine-tuned to follow user instructions and hence can better adapt to specific reasoning tasks. Table 3 presents LLM details.
| Model Class | Model Version | Size | Context Window |
| GPT-4 | gpt-4-0613 | N/A | 8k |
| GPT-3.5 | gpt-3.5-turbo-0613 | N/A | 4k |
| CodeLlama-34B | CodeLlama-34B-Instruct | 34B | 16k |
| CodeLlama-13B | CodeLlama-13B-Instruct | 13B | 16k |
| CodeLlama-7B | CodeLlama-7B-Instruct | 7B | 16k |
2.4. Prompting Strategies for LLMs
We explore various prompting strategies that can assist LLMs in predicting if a given code snippet is vulnerable. The LLMs discussed in this study support chat interactions with two major types of prompts: the system prompt and the user prompt. The system prompt can be used to set the context for the entire conversation while user prompts can be used to provide specific details throughout the chat session. We include a system prompt at the start of each input to describe the task and expected structure of the response. Since persona assignment has been shown to improve the performance of GPT-4 on specialized tasks (Salewski et al., 2024), we add the line “You are a security researcher, expert in detecting security vulnerabilities” at the start of every system prompt to assign a persona of a Security Researcher to the model. The system prompt for all experiments ends with the statement “Provide response only in the following format:” followed by an expected structure of the response from the model. The system prompt is followed by a user prompt that varies across the various prompting strategies. In all our experiments, we incorporate the target code snippet into the user prompt without any changes.
We construct different prompting strategies to study the reasoning abilities of LLMs:
2.4.1. Basic prompt
We design a very simple prompt (shown in Listing 4 in the appendix) with the goal of understanding if the model can take a target code snippet as input and detect if it is vulnerable and determine the correct CWE as well. The prompt begins with the message “Is the following code snippet prone to any security vulnerability?” followed by the code snippet.
| OWASP | Juliet | Juliet | CVEFixes | CVEFixes | Total | |
| C/C++ | Java | C/C++ | Java | |||
| Original | 2740 | 128,198 | 56,162 | 19,576 | 3926 | 210,602 |
| Filtering | 2740 | 81,280 | 35,940 | 19,576 | 3926 | 144,002 |
| Top 25 CWE | 1478 | 11,766 | 8,506 | 12,062 | 1810 | 23,560 |
| Random Selection | 1000 | 1000 | 1000 | 1000 | 1000 | 5000 |
2.4.2. CWE specific prompt
The CWE specific prompt is presented in Listing 5 in the appendix. This prompt is similar to the Basic prompt except that it asks the model to predict if the given code snippet is vulnerable to a specific target CWE. Hence, the user prompt starts with “Is the following code snippet prone to ¡CWE¿?” followed by the code snippet. The <CWE> placeholder here contains both the ID of the CWE and the name. For instance, for CWE-22, the user prompt would start with “Is the following code snippet prone to CWE-22 (Improper Limitation of a Pathname to a Restricted Directory (Path Traversal))?” followed by the target code snippet.
2.4.3. Dataflow analysis-based prompt
In addition to the straightforward one-step analysis in the previous two prompts, we also study whether providing specific step-by-step analysis instructions can help the LLMs do better reasoning over code and make better predictions. Dataflow analysis is used by several static analysis tools to infer if there exists an unsanitized path from a source to a target node. Further, prior literature has shown step-by-step instructions can often elicit better reasoning from LLMs (Wei et al., 2022b). Motivated by these observations, we designed the CWE-DF prompt (shown in Listing 6 in appendix) that prompts the model to simulate a source-sink-sanitizer-based dataflow analysis on the target code snippet before predicting if it is vulnerable. Naturally, compared to the other prompts, this prompt generates many more tokens and is hence more costly. We provide the full prompts in Appendix A.2.
2.4.4. Other prompting strategies
We also tried other prompting strategies such as Few-shot prompting and Chain-of-thought prompting. In the few-shot prompting setup, we include two examples of the task (one with a vulnerability and one without) in the CWE specific prompt before providing the target code snippet. Few-shot prompting reported poorer results than the base CWE specific prompt while requiring more tokens. Our analysis of the few-shot prompts suggests that providing more examples might not be a useful strategy for vulnerability detection. It might be more useful to use prompts that instead elicit reasoning or explanations of some kind before detecting if the given snippet is vulnerable. With Chain-of-thought prompting, we explicitly ask the model to provide a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt. This setup did not yield better results than the Dataflow analysis-based prompt. Moreover, the reasoning chains obtained by Chain-of-thought prompting (both zero-shot and few-shot) were not as elaborate as those from the Dataflow analysis-based prompt thus limiting the ease of debugging. We provide the details of this experiment in Appendix A.3.
2.5. Dataset Processing and Selection
We perform a data processing and cleaning step for each dataset before evaluating them with LLMs. We remove or anonymize information such as commits, benchmark IDs, or vulnerability names that may provide obvious hints about the vulnerability. We skip benchmarks that are spread across multiple files, due to limitations of prompt size. Table 4 presents the details of our selection stages.
Due to the prohibitively large cost of running pre-trained LLMs, we select a subset of samples from the original datasets. We select samples corresponding to vulnerability types (or CWEs) listed in MITRE’s Top 25 Most Dangerous Software Weaknesses (MITRE Top 25 CWEs, 2023). Due to prompt size limitations, we filter out code snippets larger than 2048 tokens. Finally, because the datasets are quite large, we further randomly select 500 vulnerable and 500 non-vulnerable samples per dataset. Finally, we end up with 5000 samples across five datasets. We provide more details for each dataset in Appendix A.1.
2.6. Experimental Setup
Experiments with GPT-3.5 and GPT-4. We use the OpenAI public API to perform the experiments with GPT-3.5 and GPT-4. We use the ChatCompletions API endpoint to query the models with the prompts discussed in Section 2.4. We set the sampling temperature to 0 for obtaining deterministic predictions, the maximum number of tokens to 1024, and use the default values for all other parameters. In all our experiments, we use the top-1 prediction from the models.
Experiments with CodeLlama. We run all CodeLlama experiments on two sets of machines: one with Intel Xeon machine, with 40 CPUs, four GeForce RTX 2080 Ti GPUs, and 750GB RAM, and another Intel Xeon machine with 48 CPUs, 8 A100s, and 1.5T RAM. Similar to GPT-4, we set the temperature to 0, the maximum tokens to 1024, and use top-1 prediction for evaluation.
3. Results
3.1. RQ1: Effectiveness of LLMs
We evaluate the performance of pre-trained LLMs on five open-source datasets discussed in Section 2.1. presents the best accuracy and F1 scores (across prompts) of GPT-4, GPT-3.5, and CodeLlama models (CL-7B, CL-13B, CL-34B) on all datasets. The more detailed metrics for all prompts are presented in Appendix A.4.
Modest Vulnerability Detection Performance Across LLMs. The average of the highest accuracies across all datasets is 60.6% with the maximum being 77% for CL-7B (with Basic prompt) on the Juliet Java dataset. In terms of F1 score, GPT-4 (with CWE-DF) performs the best on the synthetic datasets: OWASP (F1: 0.69), Juliet Java (F1: 0.76), and Juliet C/C++ (F1: 0.70). We observe that CodeLlama models, despite being much smaller than GPT-4, are competitive on most datasets, even with the Basic prompt. CL-34B reports the highest F1 scores (F1: 0.67) on the real-world datasets, CVEFixes Java and CVEFixes C/C++, with the CWE-DF prompt.
Performance does not improve with scale. While GPT-4 performs strictly better than GPT-3.5 on 3 of the 5 datasets, the CodeLlama models roughly exhibit similar performance. Surprisingly, CL-7B performs better than the larger LLMs on the Juliet Java and Juliet C/C++ datasets even with the Basic prompt. Listing LABEL:lst:juliet1724 presents a representative example where only CL-7B correctly predicts that an integer overflow vulnerability (CWE-190) cannot occur in the given context. Since the data variable can only contain value 2, adding 1 to it in Line LABEL:lst:line:intoverflow:add cannot cause an integer overflow. CL-7B recognizes this and predicts that the snippet is not vulnerable. CL-13B and GPT-4 (CWE-DF) ignore the hard-coded values and predict that the snippet is vulnerable, indicating that smaller LLMs may sometimes perform better than larger models.
[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]javacode/int_overflow_juliet_java_c13wrong.txt
3.2. RQ2: Comparison of Prompting Strategies
Figure 2 presents the accuracy and F1 scores (averaged across all LLMs) of the three prompting strategies across all datasets. Overall, the three prompts perform similarly in terms of accuracy. CWE-DF performs better than CWE and Basic prompt in terms of F1 score on the real-world datasets (CVEFixes C/C++ and CVEFixes Java).
We next highlight qualitative differences between various prompts:
The Basic prompt often detects incorrect vulnerability classes. Table 5 presents the percentage of samples predicted as vulnerable with the Basic prompt where the predicted CWE is correct, averaged over all datasets with the same programming language. We can observe that all models predict an incorrect vulnerability in ¿ 60% and ¿ 53% of all Java and C/C++ samples predicted as vulnerable respectively. This suggests that the Basic prompt detects incorrect CWEs in roughly half of the cases.
| Language (Avg.) | GPT-4 | GPT-3.5 | CL-34B | CL-13B | CL-7B |
| Java | 0.41 | 0.34 | 0.37 | 0.38 | 0.39 |
| C/C++ | 0.29 | 0.31 | 0.33 | 0.35 | 0.47 |
Specifying the CWE in the prompt reduces false alarms. In 16 out of the 25 model-dataset combinations, the CWE specific prompt improves or retains both the accuracy and F1 score over the Basic prompt. GPT-4 with the CWE specific prompt on Juliet Java reports a 13% higher accuracy and a 0.1 higher F1 score than those with the Basic prompt. We manually inspect 10 vulnerable and 10 non-vulnerable samples from this dataset where the CWE specific prompt is correct and the Basic prompt is not. We find that not including the CWE in the Basic prompt results in predictions discussing other incorrect CWEs. The Basic prompt incorrectly predicts that the 10 non-vulnerable samples are vulnerable to other CWEs of which we only find 3 to be plausible but unlikely due to the context. For example, it predicts that a value read from an input stream can be vulnerable if not validated (CWE-20) but this value is not used in a vulnerable context. Of the 10 vulnerable samples, Basic prompt predicts a plausible CWE for one sample while incorrectly predicting that the remaining 9 samples are not vulnerable. On the other hand, the CWE specific prompt also provides useful high-level explanations for why the snippet is vulnerable/not vulnerable in 18 / 20 samples. The 2 incorrect explanations are artifacts of faulty reasoning or hallucination: for example, an Integer Overflow due to addition to INT_MAX in the function is incorrectly attributed to subtracting from INT_MIN in the explanation.
Dataflow analysis identifies CWE-relevant textual cues and provides more elaborate explanations. The Dataflow analysis-based prompt (CWE-DF) prompt performs better than CWE specific prompt both in terms of accuracy and F1 score in 12 out of 25 model-dataset combinations and reports a higher F1 score in 23 out of 25 cases. The most significant improvement with the CWE-DF prompt over the CWE specific prompt is for the CVEFixes Java dataset and GPT-3.5 with 7% higher accuracy and 0.4 higher F1 score. We manually inspect 10 vulnerable and 10 non-vulnerable samples from this dataset where only the CWE-DF prompt is correct. We find that the CWE-DF prompt leverages textual cues for sanitization (e.g., csrftokenhandler() suggests protection from CSRF) while the CWE specific prompt does not. The CWE-DF prompt also assumes that any external parameter or call source is tainted by design and predicts that the sample is vulnerable more often (as is evident from higher F1). Further, the analysis from CWE-DF prompt is more detailed and useful in localization than the explanations from CWE specific prompt.
[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_51200_highlighted.txt
3.3. RQ3: Performance of LLMs across CWEs
We next evaluate how the LLMs perform on different classes of security vulnerabilities (CWEs). For each dataset and model, we consider the best-performing prompt for the analysis and only report CWEs with at least 10 samples. Because the CWE-wise distribution of vulnerable and non-vulnerable samples can be imbalanced, we compute balanced accuracy for each CWE (for ease of presentation, we refer to it as accuracy henceforth in this section). Figure LABEL:fig:java_cwe_distribution presents the CWE-wise distribution of accuracies on the OWASP, Juliet Java, and CVEFixes Java datasets. Figure 3 reports the accuracies on the Juliet C/C++ and CVEFixes C/C++ datasets respectively. For each model, we consider the prompt with the best accuracy on the dataset and break ties using the F1 score.
LLMs perform well on vulnerabilities that do not require additional context. We find that there are certain CWEs where multiple LLMs perform better on: Out-of-bounds Read / Write (CWE-125, CWE-787), Null pointer dereference (CWE-476) and Integer Overflow (CWE-190). The higher performance on these vulnerabilities can be attributed to the fact that these are fairly self-contained and little additional context is needed to detect them. Concretely, GPT-3.5 consistently performs well on NULL Pointer Dereference across all datasets with accuracies of 80.8%, 87.9% and 60.5% on the Juliet C/C++, Juliet Java and CVEFixes C/C++ datasets respectively. GPT-4 also performs consistently better on NULL Pointer Dereference and Integer Overflow on the synthetic datasets with accuracies of 62.1% and 68.4% on Juliet C/C++ and 78.5% and 73% on Juliet Java respectively. CL-7B reports high accuracies on Integer Overflow on the synthetic datasets with accuracies of 81% on Juliet Java and 70% on Juliet C/C++ respectively. GPT-4 and CL-34B perform extremely well on Out-of-bounds Read (78.6% and 64.3%) and Out-of-bounds Write (78.1% and 71.9%) on the CVEFixes Java dataset.
Poor performance on real-world C/C++ is due to missing global context. We see that the performance of all LLMs on vulnerabilities in CVEFixes C/C++ is worse than that on the same CWEs in CVEFixes Java and Juliet C/C++. For instance, while GPT-4 and CL-34B perform extremely well on the Out-of-bounds Read / Write vulnerabilities in CVEFixes Java as discussed above, they report accuracies ¡ 53% for these CWEs on the CVEFixes C/C++ dataset. We attribute this disparity to the nature of these vulnerabilities in the two languages: Out-of-bounds Reads / Writes in CVEFixes C/C++ require reasoning about pointers and structs which requires more context about the structs and their members. In CVEFixes Java, on the other hand, these vulnerabilities arise primarily due to illegal array indexing. This issue does not emerge in Juliet C/C++ because all the information about the pointers is presented in the snippet. We present examples in Appendix A.8.
Some LLMs are better at detecting certain CWEs. GPT-4 reports high accuracies on CWE-78 (OS Command Injection) in 3/5 datasets (CVEFixes C/C++ with 70.8%, CVEFixes Java with 61.7%, and Juliet Java with 63.8%) but the other models do not. Among the CodeLlama models, CL-13B performs well on CWE-362 (Race Condition) in CVEFixes C/C++ with 63.1% and CVEFixes Java with 66.7%. CWE-89 (SQL Injection) is another vulnerability where only CL-7B reports high accuracies on 2/3 datasets (Juliet Java with 74% and CVEFixes Java with 64%).
3.4. RQ4: Impact of Adversarial Attacks
Experiment setup. We implement three semantics-preserving attacks, proposed by Gao et al. (Gao et al., 2023): dead-code injection, variable renaming, and branch insertion (illustrated in Table 11 in appendix). We only select test samples where the original prediction of LLMs (with the CWE prompt) was correct. We select 100 test samples per dataset for this experiment and measure the reduction in accuracy after the attack. For this experiment, we select two models, GPT-4 and CL-7B, and four datasets. Figure 4 presents our results.
Results. We find GPT-4 suffers mild degradation in accuracy (average 8.6%) across all datasets and attacks. For GPT-4, the degradation is more significant for real-world datasets (up to 23%) compared to synthetic ones (up to only 2%). In contrast, for CodeLlama-7B, the accuracy significantly reduces by almost 40-50% in many cases and by 39% on average, especially for C++.
Among the three attacks, branch insertion (27% avg. reduction) and variable renaming (23% avg. reduction) have the highest impact on LLM performance. On further analysis of incorrect predictions, we observe that under these attacks, GPT-4 either fails to infer any sources, sinks, or sanitizers (even if they exist) or the reasoning chain breaks (which worked previously). Interestingly, under these attacks, CodeLlama-7B fails to follow the instructions in the system prompt and does not perform proper analysis using the dataflow analysis steps, causing it to output an incorrect response.
3.5. RQ5: LLMs vs Static Analysis Tools
Experiment setup. We next explore how GPT-4 (CWE-DF) compares against CodeQL. Since CodeQL requires building projects before analysis and the real-world datasets contain large projects, we limit our focus to the three synthetic datasets, namely OWASP and Juliet Java and Juliet C/C++. In this study, we run the official CodeQL queries designed for the top 25 CWEs. Table 6 presents results from CodeQL and GPT-4 on the three datasets. CodeQL performs better than GPT-4 (CWE-DF) in terms of accuracy on all three datasets. GPT-4 reports a 0.09 higher F1 than CodeQL on Juliet C/C++ while CodeQL reports a 0.03 and 0.18 higher F1 on OWASP and Juliet Java respectively.
| Dataset/Model | CodeQL | GPT-4 (CWE-DF) | ||||||
| A | P | R | F1 | A | P | R | F1 | |
| OWASP | 0.63 | 0.58 | 0.96 | 0.72 | 0.55 | 0.53 | 1.00 | 0.69 |
| Juliet Java | 0.94 | 0.92 | 0.95 | 0.94 | 0.70 | 0.63 | 0.98 | 0.76 |
| Juliet C/C++ | 0.72 | 0.99 | 0.44 | 0.61 | 0.59 | 0.55 | 0.98 | 0.70 |
[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_43583_highlighted.txt
3.6. RQ6: LLMs vs Deep-Learning-Based Tools
We compare LLMs against two prior deep learning-based approaches: 1) DeepDFA (Steenhoek et al., 2024), which trains Graph Neural Networks using embeddings of control flow graphs and associated data flow facts, and 2) LineVul (Fu and Tantithamthavorn, 2022), which is a transformed-based model trained using token-based representation of code.
Experiment setup. We used the DeepDFA and LineVul versions from DeepDFA’s latest artifact version (DeepDFA artifact, 2024). For CVEFixes C/C++ and Juliet C/C++, we use the same test set as our main evaluation §3.1. We split the remaining dataset into training (80%) and validation (20%) sets. We also compare the results with the BigVul (Fan et al., 2020) dataset, used in the prior work. Because BigVul is a much larger real-world dataset, it can potentially improve the performance of the tools. We explore two evaluation settings: 1) train and test on the same dataset and 2) train on one and test on a different dataset to evaluate the generalizability of these models. Table 7 presents the results. Each cell reports the average results across three runs.
DL-based approaches have limited effectiveness on real-world datasets. We observe that both DeepDFA and LineVul obtain a maximum F1 score of 0.62 on CVEFixes C/C++, which is quite similar to the scores of GPT-4. For DeepDFA, while training on larger BigVul and Juliet C/C++ datasets improves its F1 scores, the accuracy scores remain close to 50%. On the other hand, LineVul is able to learn better even when using CVEFixes C/C++, which is relatively smaller. Interestingly, when trained on BigVul, LineVul obtains a poor F1 score of 0.02 on CVEFixes C/C++, which might indicate an over-fitting problem. We observe a similar pattern when LineVul is tested on Juliet C/C++ using CVEFixes C/C++ or BigVul training datasets. For Juliet C/C++, both tools obtain better scores compared to GPT-4. LineVul obtains a perfect score indicating that Juliet C/C++ has limited code patterns that are likely easier to learn. However, as we explain later, LineVul does not generalize well.
Generalizability across datasets. We observe that, for LineVul, the F1 scores drop drastically when trained on BigVul and tested on CVEFixes C/C++ or Juliet C/C++. Compared to LineVul, DeepDFA generalizes better, obtaining higher accuracy and F1 scores on Juliet C/C++ when trained using other datasets.
Trade-offs. DeepDFA involves significant inference overhead, due to the CFG extraction and dataflow analysis steps. LLMs, however, can use the textual representation of code and can operate on incomplete/partial programs. The use of data-flow and control-flow information in DeepDFA is evidently useful. We made similar observations with LLMs when using the CWE-DF prompt. On the other hand, LineVul, like LLMs, can leverage natural language information but has a generalization problem. Finally, both DeepDFA and LineVul provide binary labels and line numbers that are difficult to interpret. LLMs can additionally provide explanations, which are useful for further debugging (as shown in prior sections).
| Model | Train/Prompt | Test | A | P | R | F1 |
| DeepDFA | BigVul | BigVul | 0.98 | 0.53 | 0.92 | 0.67 |
| LineVul | BigVul | BigVul | 0.99 | 0.96 | 0.88 | 0.92 |
| GPT-4 | CWE-DF | CVEFixes C/C++ | 0.52 | 0.51 | 0.76 | 0.61 |
| DeepDFA | CVEFixes C/C++ | CVEFixes C/C++ | 0.51 | 0.55 | 0.17 | 0.23 |
| DeepDFA | Juliet C/C++ | CVEFixes C/C++ | 0.53 | 0.53 | 0.65 | 0.58 |
| DeepDFA | BigVul | CVEFixes C/C++ | 0.52 | 0.52 | 0.76 | 0.62 |
| LineVul | CVEFixes C/C++ | CVEFixes C/C++ | 0.59 | 0.58 | 0.65 | 0.61 |
| LineVul | Juliet C/C++ | CVEFixes C/C++ | 0.50 | 0.50 | 0.91 | 0.64 |
| LineVul | BigVul | CVEFixes C/C++ | 0.50 | 0.63 | 0.01 | 0.02 |
| GPT-4 | CWE-DF | Juliet C/C++ | 0.59 | 0.55 | 0.98 | 0.70 |
| DeepDFA | Juliet C/C++ | Juliet C/C++ | 0.77 | 0.74 | 0.82 | 0.78 |
| DeepDFA | CVEFixes C/C++ | Juliet C/C++ | 0.64 | 0.72 | 0.46 | 0.55 |
| DeepDFA | BigVul | Juliet C/C++ | 0.74 | 0.70 | 0.82 | 0.77 |
| LineVul | Juliet C/C++ | Juliet C/C++ | 1.0 | 1.0 | 0.99 | 1.0 |
| LineVul | CVEFixes C/C++ | Juliet C/C++ | 0.51 | 0.62 | 0.14 | 0.22 |
| LineVul | BigVul | Juliet C/C++ | 0.43 | 0.42 | 0.39 | 0.41 |
4. Related Work
Static analysis tools for vulnerability detection. Static analysis tools search for pre-defined vulnerability patterns in code. Tools such as FlawFinder (FlawFinder, 2023) and CppCheck (CPPCheck, 2023) use syntactic and simple semantic analysis techniques to find vulnerabilities in C++ code. More advanced tools like CodeQL (Avgustinov et al., 2016), Infer (Fb Infer, 2023), and CodeChecker (Code Checker, 2023) employ semantic analysis techniques and can detect vulnerabilities in multiple languages. Static analysis tools rely on manually crafted rules and precise specifications of code behavior, which is difficult to obtain automatically. In contrast, while LLMs cannot always reliably perform end-to-end reasoning over code, we observe that LLMs can automatically identify such specifications by leveraging statistically learned rules from training data. Further, due to their semantic understanding of natural language, LLMs can also perform more contextual reasoning. These abilities can potentially be leveraged to improve static analysis tools.
Deep Learning-based vulnerability detection. Several works have focused on using Deep Learning techniques for vulnerability detection. Earlier works such as Devign (Zhou et al., 2019), Reveal (Chakraborty et al., 2020), LineVD (Hin et al., 2022) and IVDetect (Li et al., 2021) leveraged Graph Neural Networks (GNNs) for modeling dataflow graphs, control flow graphs, abstract syntax trees and program dependency graphs. Other works explored alternate model architectures: VulDeePecker (Li et al., 2020) and SySeVR (Li et al., 2018) used LSTM-based models on slices and data dependencies while Draper used Convolutional Neural Networks. Recent works demonstrate that Transformer-based models fine-tuned on the task of vulnerability detection can outperform specialized techniques (CodeBERT, LineVul (Fu and Tantithamthavorn, 2022), UnixCoder). DeepDFA (Steenhoek et al., 2024) and ContraFlow (Cheng et al., 2022) learn specialized embeddings that can further improve the performance of Transformer-based vulnerability detection tools. To the best of our knowledge, these techniques provide binary labels for vulnerability detection and cannot classify the type of vulnerability. Thapa et al. (2022) explore whether Language Models fine-tuned on multi-class classification can perform well where the classes correspond to groups of similar types of vulnerabilities. In contrast, we study some of the largest Language Models, such as GPT-4, and perform a much granular CWE-level classification, generate human-readable informal specifications and explore various prompting techniques that allow using the LLMs out-of-the-box.
LLMs for automated software engineering. Several recent approaches have demonstrated that LLMs can be effectively leveraged to improve the state-of-the-art performance in various traditional software engineering tasks such as automated program repair (Xia et al., 2023; Joshi et al., 2023; Xia and Zhang, 2022), test generation (Lemieux et al., 2023; Deng et al., 2023), code evolution (Zhang et al., 2023), and fault localization (Yang et al., 2024). However, unlike these approaches, we find that LLMs have limited vulnerability detection abilities.
Recently, Li et al. (Li et al., 2024) developed Llift, an approach that combines LLMs with static analysis to detect Use Before Initialization (UBI) bugs in Linux kernel. While they focus on a specific class of bugs, their approach supports the observations we make in Section 3.5 on the complementary nature of static analysis and LLMs. To the best of our knowledge, our work is the first comprehensive and general study of vulnerability detection abilities of LLMs across a broad range of vulnerabilities (25 CWEs) across two languages.
5. Threats to Validity
External. The choice of LLMs and datasets may bias our evaluation and insights. To address this threat, we choose multiple popular synthetic and real-world datasets across two languages: Java and C++. We also choose both state-of-the-art closed-source and open-source LLMs. However, our insights may not generalize to other languages or datasets not studied in this paper.
Internal. Owing to the non-deterministic nature of LLMs and single experiment runs per benchmark, our observations may be biased. To mitigate this threat, we use a temperature of zero to ensure determinism across all LLMs. While this works well for locally run CodeLlama models, it is well-known that GPT-4 and GPT-3.5 might still return non-deterministic results. However, due to the large number of benchmarks we evaluate, the non-determinism should balance out across the datasets. Further, given the poor effectiveness of all LLMs across the board, we do not expect our results to significantly change with re-runs.
Our evaluation code and scripts may have bugs, which might bias our results. Our manual analysis of results may lead to erroneous inferences. To address this threat, multiple co-authors reviewed the code regularly and actively fixed issues. Further, multiple co-authors of the paper independently analyzed the results and discussed them together to mitigate any discrepancies.
6. Conclusion
In this work, we performed a comprehensive analysis of LLMs for security vulnerability detection. Our study reveals that both closed-source LLMs, such as GPT-4, and open-source LLMs, like CodeLlama, perform modestly at vulnerability detection for both Java and C/C++. Their performance on vulnerability detection is lower when dealing with real-world code. However, we find that even in cases where the models produce incorrect predictions, they identify relevant sources, sinks and sanitizers for dataflow analysis. Hence, we believe that an interesting future direction is to develop neuro-symbolic techniques that combine the intuitive reasoning abilities of LLMs with symbolic tools such as logical reasoning engines and static code analyzers for more effective and interpretable solutions.
Appendix A Appendix
A.1. Dataset Processing and Selection
We perform a data processing and cleaning step for each dataset before evaluating them with LLMs.
OWASP. We remove or anonymize information in OWASP benchmarks that may provide obvious hints about the vulnerability in a file. For instance, we change package, variable names, and strings such as “owasp”, “testcode”, and “/sqli-06/BenchmarkTest02732” to other pre-selected un-identifying names such as “pcks”, “csdr”, etc. We remove all comments in the file because they may explicitly highlight the vulnerable line of code or may have irrelevant text (such as copyright info), which may leak information. These changes, however, do not change the semantics of the code snippets.
Juliet Java and C/C++. Similar to OWASP, we remove all comments and transform all identifiers that leak identifying information in all test cases in the Juliet benchmark. For instance, we change “class CWE80_XSS_CWE182_Servlet_connect_tcp_01” to “class MyClass”. The Juliet benchmark provides the vulnerable (named as “bad”) and non-vulnerable (named as “good*”) methods in the same file. For easier evaluation, we perform a pre-processing step to split each file into two, each containing either a vulnerable or non-vulnerable method. Juliet also contains special benchmarks that have dependencies across multiple (2-5) files. We skip these benchmarks because they are typically too big to fit into the LLM prompt. Hence, the number of test cases after the data processing step in Juliet is reduced (as shown in Table 4).
CVEFixes. For each CVE, CVEFixes provides the methods that were involved in the fix commit. It also includes the method code in the parent commit, i.e., the method version before the fix. We collect all methods in the fix commit and the parent commit and label them as vulnerable and non-vulnerable, respectively. Similar to other datasets, we also remove all comments in the method code. While CVEFixes contains methods across multiple programming languages, we only collect C/C++ and Java methods for our study.
A.2. Prompting Strategies
The Basic prompt is presented in Listing 1, CWE specific prompt in Listing 2 and Dataflow analysis-based prompt in Listing 3.
A.3. Other Prompting Strategies
In addition to the prompting strategies presented in our main evaluation, we considered other popular prompting strategies such as Few-shot prompting and Chain-of-thought prompting in a limited experimental setting. For the few-shot prompt (CWE-Few-shot), we included two examples of the task (one with a vulnerability and one without) in the CWE specific prompt before providing the target code snippet. For the chain-of-thought prompt (CWE-CoT), we explicitly ask the model to provide a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt.
Table 8 and Table 9 present the results from GPT-4 with various prompting strategies on a random subset of 100 samples of the Juliet Java and CVEFixes C/C++ datasets respectively. The CWE-DF prompt reports the highest accuracy of 69% and the highest F1 score of 0.75 on the Juliet Java dataset. The CWE-DF prompt reports a 0.05 higher F1 score than the CWE-CoT prompt and a 0.03 higher F1 score than the CWE-Few-shot prompt. This difference is much more prominent on the CVEFixes C/C++ dataset where the CWE-DF prompt reports a 0.34 higher F1 score than the CWE-CoT prompt and a 0.31 higher F1 score than the CWE-Few-shot prompt. Moreover, the CWE-Few-shot prompt reported a 0.2 lower F1 score than the CWE specific prompt on the CVEFixes C/C++ dataset while requiring more tokens. Our analysis of the few-shot prompts suggests that providing more examples may not be a useful strategy for vulnerability detection. Because the potential set of vulnerable code patterns is quite large, the provided examples hardly make a difference to LLMs’ reasoning abilities. Hence, it may be more useful to use prompts that instead elicit reasoning or explanations of some kind before detecting if the given snippet is vulnerable. The CWE-CoT prompt, however, does not help with reasoning always, as it either performed at par or worse than the Dataflow analysis-based prompt.
| Model | Prompt | Metrics | ||||
| A | P | R | F1 | |||
| GPT-4 | CWE | 0.65 | 0.58 | 0.96 | 0.72 | |
| GPT-4 | CWE-Few-shot | 0.65 | 0.58 | 0.94 | 0.72 | |
| GPT-4 | CWE-CoT | 0.69 | 0.64 | 0.79 | 0.70 | |
| GPT-4 | CWE-DF | 0.69 | 0.61 | 0.96 | 0.75 | |
| Model | Prompt | Metrics | |||||
| A | P | R | F1 | ||||
| GPT-4 | CWE | 0.55 | 0.54 | 0.58 | 0.56 | ||
| GPT-4 | CWE-Few-shot | 0.49 | 0.38 | 0.34 | 0.36 | ||
| GPT-4 | CWE-CoT | 0.52 | 0.37 | 0.30 | 0.33 | ||
| GPT-4 | CWE-DF | 0.56 | 0.56 | 0.83 | 0.67 | ||
Learning from these experiments, we selected the CWE specific prompt, Dataflow analysis-based prompt, in addition to the Basic prompt, for our main evaluation with LLMs.
A.4. Detailed metrics across all LLMs and Datasets
Table 10 presents the metrics for all LLMs and Datasets across all prompts.
| Model | Prompt | OWASP | Juliet Java | CVEFixes Java | Juliet C/C++ | CVEFixes C/C++ | ||||||||||||||||||||
| A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | |||||||
| GPT-4 | Basic | 0.52 | 0.51 | 1.00 | 0.68 | 0.56 | 0.54 | 0.85 | 0.66 | 0.50 | 0.50 | 0.34 | 0.41 | 0.54 | 0.52 | 0.92 | 0.67 | 0.51 | 0.50 | 0.57 | 0.54 | |||||
| GPT-4 | CWE | 0.54 | 0.52 | 1.00 | 0.69 | 0.69 | 0.63 | 0.97 | 0.76 | 0.55 | 0.56 | 0.44 | 0.50 | 0.58 | 0.54 | 0.95 | 0.69 | 0.52 | 0.52 | 0.52 | 0.52 | |||||
| GPT-4 | CWE-DF | 0.55 | 0.53 | 1.00 | 0.69 | 0.70 | 0.63 | 0.98 | 0.76 | 0.53 | 0.53 | 0.59 | 0.56 | 0.59 | 0.55 | 0.98 | 0.70 | 0.52 | 0.51 | 0.76 | 0.61 | |||||
| GPT-3.5 | Basic | 0.53 | 0.52 | 0.72 | 0.60 | 0.58 | 0.57 | 0.71 | 0.63 | 0.46 | 0.35 | 0.09 | 0.15 | 0.49 | 0.49 | 0.64 | 0.56 | 0.52 | 0.55 | 0.19 | 0.29 | |||||
| GPT-3.5 | CWE | 0.55 | 0.54 | 0.62 | 0.58 | 0.52 | 0.52 | 0.55 | 0.54 | 0.47 | 0.41 | 0.12 | 0.19 | 0.49 | 0.49 | 0.70 | 0.58 | 0.51 | 0.54 | 0.19 | 0.28 | |||||
| GPT-3.5 | CWE-DF | 0.51 | 0.51 | 0.93 | 0.66 | 0.40 | 0.44 | 0.73 | 0.55 | 0.54 | 0.53 | 0.66 | 0.59 | 0.40 | 0.44 | 0.77 | 0.56 | 0.52 | 0.52 | 0.75 | 0.61 | |||||
| CL-34B | Basic | 0.51 | 0.51 | 1.00 | 0.67 | 0.47 | 0.48 | 0.85 | 0.62 | 0.50 | 0.50 | 0.28 | 0.36 | 0.50 | 0.50 | 0.93 | 0.65 | 0.51 | 0.51 | 0.19 | 0.28 | |||||
| CL-34B | CWE | 0.57 | 0.54 | 0.94 | 0.69 | 0.49 | 0.49 | 0.94 | 0.65 | 0.50 | 0.51 | 0.16 | 0.25 | 0.53 | 0.52 | 0.98 | 0.68 | 0.51 | 0.54 | 0.08 | 0.14 | |||||
| CL-34B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | |||||
| CL-13B | Basic | 0.52 | 0.51 | 0.98 | 0.67 | 0.47 | 0.48 | 0.72 | 0.58 | 0.50 | 0.50 | 0.13 | 0.21 | 0.46 | 0.48 | 0.79 | 0.59 | 0.51 | 0.52 | 0.22 | 0.31 | |||||
| CL-13B | CWE | 0.52 | 0.51 | 0.98 | 0.67 | 0.50 | 0.50 | 0.89 | 0.64 | 0.48 | 0.47 | 0.29 | 0.36 | 0.53 | 0.51 | 0.98 | 0.67 | 0.52 | 0.52 | 0.56 | 0.54 | |||||
| CL-13B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 0.96 | 0.66 | |||||
| CL-7B | Basic | 0.56 | 0.54 | 0.81 | 0.65 | 0.77 | 0.79 | 0.72 | 0.75 | 0.47 | 0.46 | 0.42 | 0.44 | 0.62 | 0.58 | 0.86 | 0.69 | 0.47 | 0.35 | 0.07 | 0.11 | |||||
| CL-7B | CWE | 0.50 | 0.50 | 1.00 | 0.67 | 0.52 | 0.51 | 0.99 | 0.67 | 0.51 | 0.51 | 0.84 | 0.63 | 0.51 | 0.50 | 0.99 | 0.67 | 0.50 | 0.50 | 0.84 | 0.63 | |||||
| CL-7B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.51 | 0.50 | 0.98 | 0.66 | |||||
A.5. Adversarial Attack Examples
| Attack | Example |
| Dead-code Injection | ⬇ int obj = new Object(...); if (obj == null) { /* deadcode */ } |
| Variable Renaming | ⬇ String var3 = (String) names.nextElement(); String[] values = request.getParameterValues(var3); |
| Branch Insertion | ⬇ boolean var11 = true; if (var11) { configManager.init(); } |
Table 11 presents the three adversarial attacks that we implemented: dead-code injection, variable renaming, and branch insertion. Each attack contains multiple code patterns, which are randomly chosen when perturbing a given code snippet. We implement these attacks for Java and C/C++. For a given code snippet, we only apply an attack only once.
A.6. More Details on CodeQL vs. GPT-4
Table 12 presents the CWE-wise distribution of CodeQL results. CodeQL shows relatively better performance on Juliet Java (with an F1 score over 0.75 across CWEs) than Juliet C/C++ (with two CWEs reporting F1 scores under 0.4). On Juliet Java, CodeQL re- ports the highest performance on CWE-78 with an F1 score of 0.92. Interestingly, CWE-78 reports the worst results on the other two datasets (F1 score of 0.67 with OWASP and 0.03 with Juliet C/C++).
[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_47137_highlighted.txt
In Listing LABEL:fig:julietcpp47137, the model incorrectly reasons about the sanitization by overlooking the & which would cause both the expressions to be evaluated. The identified sources and sinks are correct, however.
| Dataset | CWE | CodeQL | GPT-4 (CWE-DF) | ||||||
| A | P | R | F1 | A | P | R | F1 | ||
| OWASP | 22 | 0.52 | 0.50 | 1.00 | 0.67 | 0.48 | 0.48 | 1.00 | 0.64 |
| OWASP | 78 | 0.56 | 0.53 | 0.77 | 0.63 | 0.48 | 0.48 | 1.00 | 0.65 |
| OWASP | 89 | 0.57 | 0.54 | 1.00 | 0.70 | 0.52 | 0.52 | 1.00 | 0.68 |
| OWASP | 79 | 0.79 | 0.71 | 1.00 | 0.83 | 0.65 | 0.59 | 1.00 | 0.74 |
| Juliet Java | 78 | 0.94 | 0.86 | 1.00 | 0.92 | 0.57 | 0.47 | 1.00 | 0.64 |
| Juliet Java | 89 | 0.88 | 0.85 | 1.00 | 0.92 | 0.69 | 0.68 | 1.00 | 0.81 |
| Juliet Java | 476 | 0.76 | 0.76 | 0.76 | 0.76 | 0.79 | 0.75 | 0.88 | 0.81 |
| Juliet Java | 190 | 0.97 | 1.00 | 0.92 | 0.96 | 0.71 | 0.62 | 0.97 | 0.75 |
| Juliet C/C++ | 78 | 0.52 | 1.00 | 0.04 | 0.08 | 0.50 | 0.50 | 1.00 | 0.67 |
| Juliet C/C++ | 476 | 0.80 | 0.95 | 0.75 | 0.84 | 0.71 | 0.75 | 0.86 | 0.80 |
| Juliet C/C++ | 416 | 0.63 | 1.00 | 0.26 | 0.41 | 0.52 | 0.51 | 0.97 | 0.67 |
| Juliet C/C++ | 190 | 0.93 | 0.99 | 0.87 | 0.93 | 0.68 | 0.60 | 0.98 | 0.74 |
A.7. Qualitative analysis of GPT-4 responses
We first present examples where the dataflow analysis from the CWE-DF prompt is useful. Consider the code snippet in Listing 4. In this snippet, the variable dir is indirectly being used to create a directory via the dirToCreate variable. GPT-4 correctly identifies that this path is not sanitized and could be used to create a directory in otherwise restricted locations. This could lead to CWE-22 (path traversal) as is rightly predicted by GPT-4. In Listing 5, the method write takes an array b and integers offset and length as inputs. The expression offset + length > b.length can lead to an Integer Overflow (CWE-190) if offset + length exceeds the maximum integer value and this is correctly identified by GPT-4. Finally, in Listing 6, there are multiple array accesses that could lead to CWE-125 (Out-of-bounds read). GPT-4 correctly identifies that the expression ciphertextOffset + length + index could exceed the size of the ciphertext array, leading to this vulnerability.
A.8. CWE-125 (Out-of-bounds Read in Java vs. C++)
Listing 6 presents an example of Out-of-bounds read vulnerability in CVEFixes Java which GPT-4 predicts correctly. Codellama-34B also correctly predicts that this snippet is vulnerable. Listing 7 presents an example from CVEFixes C/C++ where Out-of-bounds read vulnerability is not possible but this is not evident from context. Codellama-34B’s explanation seems plausible given that little is known about the pointers in question.

### v3

@@
 Understanding the Effectiveness of Large Language Models in Detecting Security Vulnerabilities
-Abstract.
+Abstract
 Security vulnerabilities in modern software are prevalent and harmful.
-While automated vulnerability detection tools have made promising progress, their scalability and applicability remain challenging.
-Recently, Large Language Models (LLMs), such as GPT-4 and CodeLlama, have demonstrated remarkable performance on code-related tasks.
-However, it is unknown whether such LLMs can do complex reasoning over code.
-In this work, we explore whether pre-trained LLMs can detect security vulnerabilities and address the limitations of existing tools.
-We evaluate the effectiveness of pre-trained LLMs, in terms of performance, explainability, and robustness, on a set of five diverse security benchmarks spanning two languages, Java and C/C++, and covering both synthetic and real-world projects.
-Overall, all LLMs show modest effectiveness in end-to-end reasoning about vulnerabilities, obtaining an average of 60% accuracy across all datasets.
-However, we observe that LLMs show promising abilities at performing parts of the analysis correctly, such as identifying vulnerability-related specifications (e.g., sources and sinks) and leveraging natural language information to understand code behavior (e.g., to check if code is sanitized).
-Further, LLMs are relatively much better at detecting simpler vulnerabilities that typically only need local reasoning (e.g., Integer Overflows and NULL pointer dereference).
-We find that advanced prompting strategies that involve step-by-step analysis significantly improve performance of LLMs on real-world datasets (improving F1 score by up to 0.25 on average).
-Finally, we share our insights and recommendations for future work on leveraging LLMs for vulnerability detection.
-1. Introduction
+While automated vulnerability detection techniques have made promising progress, their scalability and applicability remain challenging.
+The remarkable performance of Large Language Models (LLMs), such as GPT-4 and CodeLlama, on code-related tasks has prompted recent works to explore if LLMs can be used to detect security vulnerabilities.
+In this paper, we perform a more comprehensive study by examining a larger and more diverse set of datasets, languages, and LLMs, and qualitatively evaluating detection performance across prompts and vulnerability classes.
+Concretely, we evaluate the effectiveness of 16 pre-trained LLMs on 5,000 code samples—1,000 randomly selected each from five diverse security datasets.
+These balanced datasets encompass synthetic and real-world projects in Java and C/C++ and cover 25 distinct vulnerability classes.
+Our results show that LLMs across all scales and families show modest effectiveness in end-to-end reasoning about vulnerabilities, obtaining an average accuracy of 62.8% and F1 score of 0.71 across all datasets.
+LLMs are significantly better at detecting vulnerabilities that typically only need intra-procedural reasoning, such as OS Command Injection and NULL Pointer Dereference.
+Moreover, LLMs report higher accuracies on these vulnerabilities than popular static analysis tools, such as CodeQL.
+We find that advanced prompting strategies that involve step-by-step analysis significantly improve performance of LLMs on real-world datasets in terms of F1 score (by up to 0.18 on average).
+Interestingly, we observe that LLMs show promising abilities at performing parts of the analysis correctly, such as identifying vulnerability-related specifications (e.g., sources and sinks) and leveraging natural language information to understand code behavior (e.g., to check if code is sanitized).
+We believe our insights can motivate future work on LLM-augmented vulnerability detection systems.
+I Introduction
 Security vulnerabilities afflict software despite decades of advances in programming languages, program analysis tools, and software engineering practices.
-Even well-tested and critical software such as OpenSSL, a widely used library for applications that provide secure communications, contains trivial buffer overflow vulnerabilities, e.g., (CVE-2022-3602, 2022) and (CVE-2022-3786, 2022).
-A recent study by Microsoft showed that more than 70% of the vulnerabilities are still caused by well-understood memory safety issues (Miller, 2019).
-The size and complexity of modern software systems are growing quickly, encompassing numerous programs, libraries, and modules that interact with each other.
+Even well-tested and critical software such as OpenSSL, a widely used library for applications that provide secure communications, contains trivial buffer overflow vulnerabilities, e.g., [1] and [2].
+A recent study by Microsoft showed that more than 70% of vulnerabilities are still caused by well-understood memory safety issues [3].
+This is alarming given the rapidly growing size and complexity of modern software systems, encompassing numerous programs, libraries, and modules that interact with each other.
 Hence, we need major technical advances to effectively detect security vulnerabilities in such complex software.
-Traditional techniques for automated vulnerability detection, such as fuzzers (Manes et al., 2018), and static analyzers such as CodeQL (Avgustinov et al., 2016) and Semgrep (Semgrep, 2023) have made promising strides.
-For example, in the last two years, researchers found over 300 security vulnerabilities through custom CodeQL queries (Semmle, 2023; Leong, 2022).
+Traditional techniques for automated vulnerability detection, such as fuzzers [4], and static analyzers such as CodeQL [5] and Semgrep [6] have made promising strides.
+For example, in the last two years, researchers found over 300 security vulnerabilities through custom CodeQL queries [7, 8].
 However, these techniques face challenges in scalability and applicability.
-Fuzzing does not scale to large applications, and fuzzing parts of applications requires manually crafting fuzz drivers.
-Moreover, it is hard to use on large critical programs with complex inputs, such as network servers, GUI-based programs, embedded firmware, boot loaders, and system services.
+Fuzzing requires manually crafted fuzz drivers and does not scale to large critical programs with complex inputs, such as network servers, embedded firmware, and system services.
 On the other hand, static analysis relies heavily on manual API specifications, and skillfully crafted heuristics to balance precision and scalability.
-In light of these challenges, GitHub pays a bounty of over 7K USD for each CodeQL query that can find new critical security bugs (GitHub, 2023).
+Until recently, GitHub paid a bounty of over 7K USD for each CodeQL query that found new critical security bugs [9].
+| Research Questions | Findings | |
+| RQ1: How do different pre-trained LLMs perform in detecting security vulnerabilities across different languages and datasets?
+(Section §III-A) | ✓ LLMs across all sizes report a mean accuracy of about 62.8% and a mean F1 score of 0.71 across all datasets.
+✗ Average accuracy on real-world datasets is 10.5% lower than that on synthetic datasets.
+✓ In stark contrast to other domains, smaller models such as Qwen-2.5-14B and Qwen-2.5-32B report higher accuracies on the real-world datasets than much larger models such as GPT-4.
+| |
+| RQ2: How do different prompting strategies affect the performance of LLMs?
+(Section III-B) | ✓ Using prompts that focus on detecting specific CWEs improves the performance of LLMs.
+✓ The Dataflow analysis-based prompt prompt further improves results for larger LLMs with an increase of up to 0.18 in F1 score on real-world datasets.
+✓ We also observe that LLMs often infer the correct sources/sinks/sanitizers but fail in end-to-end reasoning.
+| |
+| RQ3: How does the performance of LLMs vary across different vulnerability classes?
+(Section III-C) | ✓ LLMs are better at detecting local vulnerabilities that require no global context across datasets (OS Command Injection, NULL Pointer Dereference, Out-of-bounds Read/Write, etc.).
+✗ LLMs struggle to detect vulnerabilities that require additional context or reasoning about complex data structures (Out-of-bounds Read/Write with C++ structs and pointers).
+✓ Certain LLMs are very good at detecting specific CWEs across datasets (Llama-3.1-70B on OS Command Injection, DeepSeekCoder-7B on NULL Pointer Dereference, etc.).
+| |
+| RQ4: How do LLMs compare to state-of-the-art static analysis tools?
+(Section III-D) | ✗ LLMs report lower overall accuracies than CodeQL on all synthetic datasets.
+✓ LLMs report higher accuracies than CodeQL on certain vulnerability classes across datasets (Path Traversal, OS Command Injection, etc).
+CodeQL reports higher accuracies on Integer Overflow across datasets.
+| |
+| RQ5: How do LLMs compare to state-of-the-art deep-learning-based tools?
+(Section III-E) | ✓ Deep Learning(DL)-based tools such as DeepDFA [10] and LineVul [11] report accuracies similar to Qwen-2.5-32B on CVEFixes C/C++ even after being trained on in-distribution samples whereas Qwen-2.5-32B reports higher F1 scores.
+✓ DL-based tools report lower accuracies and F1 scores than LLMs when trained and evaluated on different datasets.
+✓ LLMs provide natural language explanations for their predictions while DL-based tools provide binary scores and line numbers that are often difficult to interpret.
+|
 Large Language Models (LLMs), including pre-trained models such as GPT-4 and CodeLlama, have made remarkable advances in code-related tasks in a relatively short period.
-Such tasks include code completion (Chen et al., 2021), automated program repair (Xia et al., 2023; Joshi et al., 2023; Xia and Zhang, 2022), test generation (Lemieux et al., 2023; Deng et al., 2023), code evolution (Zhang et al., 2023), and fault localization (Yang et al., 2024).
+Such tasks include code completion [12], automated program repair [13, 14, 15], test generation [16, 17], code evolution [18], and fault localization [19].
 These results clearly show the promise of LLMs, opening up a new direction for exploring advanced techniques.
@@
 This is especially important in light of the rapidly evolving landscape of LLMs in three aspects: scale, diversity, and applicability.
-First, scaling these models to ever larger numbers of parameters has led to significant improvements over previous generations in their capabilities—a phenomenon termed as emergent behavior (Wei et al., 2022a).
-For instance, GPT-4, which is presumably orders of magnitude larger than its 175-billion predecessor GPT-3.5, significantly outperforms GPT-3.5 on a wide range of code-understanding tasks (Bubeck et al., 2023).
-Second, the diversity of LLMs has grown rapidly and now includes not only proprietary general-purpose ones such as GPT-4 but also open-sourced LLMs such as CodeLlama (Rozière et al., 2023) and StarCoder (Li et al., 2023) that are specialized for code.
-Finally, the reasoning capabilities of LLMs (and hence their applicability) may vary significantly across different programming languages.
+First, scaling these models to ever larger numbers of parameters has led to significant improvements in their capabilities[20].
+For instance, GPT-4, which is presumably orders of magnitude larger than its 175-billion predecessor GPT-3.5, significantly outperforms GPT-3.5 on a wide range of code-understanding tasks [21].
+Second, the diversity of LLMs has grown rapidly and now includes not only proprietary general-purpose ones such as GPT-4 but also open-source code-specific LLMs such as CodeLlama [22] and StarCoder [23].
+Finally, the reasoning capabilities of LLMs (and hence their applicability) may vary significantly across different prompting strategies and programming languages.
 All these factors open up a large exploration space for applying LLMs to the challenging task of vulnerability detection.
+| Study Features | [24] | [25] | [26] | [27] | [28] | Our Work |
+| Languages | C/C++ | C/C++ | C/C++ | C,Py | C/C++ | C/C++,Java |
+| #Samples | 368 | 347 | 100 | 96 | 25.9 | 5000 |
+| #CWEs | 25 | N/A | N/A | 8 | 140 | 25 |
+| #LLMs (1B parameters) | 3 | 16 | 11 | 5 | 4 | 16 |
+| Comparison of various LLMs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
+| Qualitative prompt analysis | ✗ | ✗ | ✗ | ✓ | ✗ | ✓ |
+| CWE-wise analysis | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ |
+| Comparison with SA tools | ✗ | ✓ | ✗ | ✗ | ✗ | ✓ |
+| Comparison with DL tools | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ |
 Our Work.
-We conduct the first comprehensive study of using LLMs for detecting security vulnerabilities.
-We study five state-of-the-art LLMs, including proprietary models such as GPT-3.5 and GPT-4, and open-source models like CodeLlama.
-We evaluate these models on five popular security vulnerability datasets.
-We design a set of three prompting strategies for LLMs to elicit increasingly sophisticated forms of reasoning and explanations.
+We study the vulnerability detection capabilities of 16 state-of-the-art LLMs across different scales and families, including proprietary models such as Gemini and GPT-4, and open-source models like CodeLlama and Qwen.
+We evaluate these models on five popular security vulnerability datasets across two languages, C/C++ and Java, and 25 vulnerability classes.
+We first study how LLMs perform on the task of vulnerability detection using three prompting strategies and how these strategies qualitatively compare against each other.
+We also attempt to identify vulnerability classes that most benefit from the use of LLMs and these prompting techniques.
 Our simplest prompting strategies include the Basic prompt, which simply asks an LLM to check for any vulnerabilities in the given code and the CWE specific prompt, which asks the LLM to check for a specific class of vulnerabilities or CWEs (such as Buffer Overflows).
-A significant limitation of static vulnerability detection tools is the requirement of building the target project to enable them to find bugs.
-Further, they also require concrete specifications of APIs (e.g., sources, sanitizers, and sinks).
-In contrast, LLMs have an internal model of APIs already seen during the pre-training phase and hence they do not require compiled or complete codebases to run.
-Inspired by this insight, we additionally design a new prompting strategy, called Dataflow analysis-based prompt, that simulates a source-sink-sanitizer based dataflow analysis on the target code snippet before predicting if it is vulnerable.
-The Dataflow analysis-based prompt, similar to a classical dataflow-based static analyzer asks the LLM to first infer the sources, sinks, and sanitizers in the code snippet and check for any unsanitized data flows between the source and sink.
-This style of prompting is similar to chain-of-thought reasoning and simulates step-by-step reasoning by LLMs.
-A key challenge in evaluating LLMs is data leakage: LLMs may perform well on a dataset because such samples were already present in their pre-training data.
-Hence, we implement three semantics-preserving adversarial attacks for code and evaluate whether they significantly impact the performance of LLMs.
-Research Questions. We study the following research questions:
-RQ1: How do different pre-trained LLMs perform in detecting security vulnerabilities across different languages and datasets? (Section §3.1)
-RQ2: How do different prompting strategies affect the performance of LLMs? (Section 3.2)
-RQ3: How does the performance of LLMs vary across different vulnerability classes? (Section 3.3)
-RQ4: Can adversarial attacks impact the performance of LLMs in detecting vulnerabilities? (Section §3.4)
-RQ5: How do LLMs compare to state-of-the-art static analysis tools? (Section 3.5)
-RQ6: How do LLMs compare to state-of-the-art deep-learning-based tools? (Section 3.6)
-Results and Findings.
-We choose five state-of-the-art LLMs: GPT-4, GPT-3.5, CodeLlama-34B, CodeLlama-13B, and CodeLlama-7B and evaluate their effectiveness in detecting security vulnerabilities across five vulnerability datasets: OWASP (OWASP Benchmark Suite, 2023), Juliet Java (Juliet Java, 2023), Juliet C/C++ (Juliet C/C++, 2023), CVEFixes Java (Bhandari et al., 2021), and CVEFixes C/C++ (Bhandari et al., 2021).
-These datasets cover 5000 vulnerable and non-vulnerable code samples, across 25 CWEs.
-Our findings are summarized as follows:
-LLMs across all sizes have modest vulnerability detection ability, with an average accuracy of only about 60% across all datasets.
-Using prompts that focus on detecting specific CWEs improves the performance of LLMs.
-The dataflow analysis-based prompt further improves results for larger LLMs by allowing step-by-step reasoning.
-The improvement is significant for real-world datasets (an increase of up to 0.25 F1 score).
-However, we observe that LLMs often infer the relevant source, sink, and sanitizers correctly but fail in end-to-end logical reasoning.
-LLMs are relatively better at detecting vulnerabilities that require local reasoning without complex code structures (such as Integer Overflow and NULL Pointer Dereference).
-However, they struggle to detect more complex vulnerabilities in real-world programs.
-Larger LLMs like GPT-4 show mild degradation in performance when subject to adversarial attacks on code (average drop 8.6%).
-In contrast, smaller LLMs, like CodeLlama-7B, show a significant drop in performance (average drop 39%) in the presence of perturbations.
-LLMs lag behind state-of-the-art static analysis tools like CodeQL in average accuracy (by 15%) and precision (by 26%) across datasets.
-However, LLMs like GPT-4 show some promising partial abilities, e.g., by identifying source/sink specifications or doing context-based reasoning using natural language information, which can potentially be useful for static analysis.
-We obtain mixed results comparing LLMs to prior deep-learning-based tools: while the deep-learning-based tools only slightly outperform LLMs on synthetic datasets (by 0.07 in F1 score), all approaches struggle equally on real-world datasets often producing close to 50% accuracy only (same as a random baseline).
-Contributions. To summarize, we make the following contributions in this paper:
-Empirical Study: We conduct the first large comprehensive study on how state-of-the-art LLMs perform in detecting security vulnerabilities across five datasets and two programming languages (C/C++ and Java).
-Prompting Strategies: We design three prompting strategies for LLMs, inspired by the recent advances in natural language processing and traditional program analysis techniques, that elicit different reasoning styles from LLMs and also provide human-readable explanations for their predictions.
-Robustness of LLMs: We study how the performance of LLMs is impacted by semantics-preserving adversarial attacks on code.
+Inspired by the success of static analysis tools like CodeQL that use dataflow analysis to detect vulnerabilities, we design a prompting strategy called Dataflow analysis-based prompt.
+This prompt asks the LLM to simulate a source-sink-sanitizer based dataflow analysis on the target code snippet before predicting if it is vulnerable.
+We next compare LLMs with existing vulnerability detection tools, namely static analysis-based CodeQL and two deep learning-based techniques, DeepDFA [10] and LineVul [11].
+As discussed earlier, static vulnerability detection tools are limited by the need for concrete API specifications and require compiling / building entire target projects before detection.
+Pre-LLM deep learning-based approaches such as DeepDFA [10] and LineVul [11] attempt to mitigate some of these limitations through fine-tuned neural representations of code.
+On the other hand, LLMs do not require the compilation of complete projects as they can be prompted to analyze partial code snippets.
+Moreover, they already have an internal model of APIs through pre-training and do not need to be trained on large datasets from scratch.
+We analyze the benefits and shortcomings of LLMs over CodeQL, including vulnerability classes where they outperform each other.
+We also study how they compare against the deep learning based-approaches in terms of generalization across datasets.
+Comparison with other studies.
+There are other studies that also evaluate the effectiveness of LLMs on the task of vulnerability detection [24, 25, 26, 27, 28].
+Table II presents a comparison of our work with these studies.
+We present our study as the most comprehensive on this topic for the following reasons:
+Size and diversity of the datasets: We curate a dataset of 5K samples, with equal number of vulnerable and non-vulnerable snippets.
+The only larger dataset is from [28] but only 695 of their 25.9K samples are vulnerable.
+Furthermore, our study is the first to include Java code.
+Comparison with non-LLM-based tools: Our study is the first to compare LLMs with CodeQL and specialized Deep Learning-based tools.
+Moreover, we also find vulnerability classes where LLMs outperform CodeQL and vice versa which is useful for deployment.
+Qualitative analysis of prompts and vulnerability classes: While other studies quantitatively compare prompting strategies, we also attempt to qualitatively identify the benefits of various prompt elements.
+Furthermore, we identify partial capabilities offered by some of these prompts that can be leveraged in LLM-based detection tools.
+We also identify vulnerability classes where LLMs perform well.
+Contributions. Our research questions and key findings are summarized in Table I. To summarize, we make the following contributions in this paper:
+Empirical Study: We conduct the largest comprehensive study on how state-of-the-art LLMs perform in detecting security vulnerabilities across 5000 samples from five datasets, two programming languages (C/C++ and Java) and covering 25 unique vulnerability classes.
 Comparison with other vulnerability detection tools: We contrast the performance of LLMs against popular static analysis and deep-learning-based vulnerability detection tools.
+We also identify vulnerability classes where LLMs perform better/worse than some of these tools.
+Qualitative comparison of prompting strategies: We quantitatively and qualitatively compare three prompting strategies, including a novel prompt inspired by dataflow analysis-based vulnerability detection.
 Insights: We perform a rigorous manual analysis of LLMs’ predictions and highlight vulnerability patterns that impact the performance of these models.
-2. Approach
-2.1. Datasets
-For our study, we select five diverse vulnerability datasets from two languages: C++ and Java.
-For each language, we also select both synthetic and real-world benchmarks.
-Table 1 presents the details of each dataset, such as the dataset size, programming language, number of vulnerable and non-vulnerable samples, and the number of unique CWEs.
-We describe each dataset next.
+II Approach
+II-A Datasets
+For our study, we select five diverse synthetic/real-world vulnerability datasets from two languages: C++ and Java.
+Table III presents the details of each dataset, such as the dataset size, programming language, number of vulnerable and non-vulnerable samples, and the number of unique CWEs.
+The synthetic benchmarks, OWASP and Juliet, allow for easy comparison with CodeQL and the real-world benchmarks, CVEFixes, are useful for evaluating practical utility.
+While many real-world datasets have been proposed in the literature, we selected CVEFixes because it is the only dataset that 1) contains vulnerability metadata such as CVE and CWE IDs, 2) is two-sided, i.e., contains both vulnerable and non-vulnerable code samples, and 3) covers multiple languages such as Java and C/C++.
+Table IV shows a comparison of existing real-world vulnerability datasets.
+We briefly describe each of the selected datasets next:
 | Dataset | Language | Size | Vul/Non-Vul | CWEs | |
-| OWASP(OWASP Benchmark Suite, 2023) | Java | 2740 | 1415/1325 | 11 | |
-| SARD Juliet (C/C++) (Juliet C/C++, 2023) | C/C++ | 81,280 | 40,640/40,640 | 118 | |
-| SARD Juliet (Java) (Juliet Java, 2023) | Java | 35,940 | 17,970/17,970 | 112 | |
-| CVEFixes (Bhandari et al., 2021) | C/C++ | 19,576 | 8223/11,347 | 131 | |
-| CVEFixes (Bhandari et al., 2021) | Java | 3926 | 1461/2465 | 68 |
-2.1.1. OWASP (Synthetic)
-The Open Web Application Security Project (OWASP) benchmark (OWASP Benchmark Suite, 2023) is a Java test suite designed to evaluate the effectiveness of vulnerability detection tools.
+| OWASP[29] | Java | 2740 | 1415/1325 | 11 | |
+| SARD Juliet (C/C++) [30] | C/C++ | 81,280 | 40,640/40,640 | 118 | |
+| SARD Juliet (Java) [31] | Java | 35,940 | 17,970/17,970 | 112 | |
+| CVEFixes [32] | C/C++ | 19,576 | 8223/11,347 | 131 | |
+| CVEFixes [32] | Java | 3926 | 1461/2465 | 68 |
+II-A1 OWASP (Synthetic)
+The Open Web Application Security Project (OWASP) benchmark [29] is a Java test suite designed to evaluate the effectiveness of vulnerability detection tools.
 Each test represents a synthetically designed code snippet containing a security vulnerability.
-OWASP contains 2740 test cases representing 11 unique classes of security vulnerabilities (also known as Common Weakness Enumeration or CWE).
-2.1.2. Juliet (Synthetic)
-Juliet (Boland and Black, 2012) is a widely-used vulnerability dataset developed by NIST.
-Juliet comprises thousands of synthetically generated test cases representing various known vulnerability patterns.
-It contains 81,280 C/C++ programs covering 118 unique CWEs, and 35,940 Java programs covering 112 unique CWEs.
-For our paper, we use the latest version, Juliet 1.3 (Black and Black, 2018; Juliet C/C++, 2023; Juliet Java, 2023).
-2.1.3. CVEFixes (Real-World)
+II-A2 Juliet (Synthetic)
+Juliet [33] is a widely-used vulnerability dataset developed by NIST with thousands of synthetically generated test cases from various known vulnerability patterns.
+II-A3 CVEFixes (Real-World)
 Bhandari et al.
-(Bhandari et al., 2021) curated a dataset, known as CVEFixes, from 5365 Common Vulnerabilities and Exposures (CVE) records from the National Vulnerability Database (NVD).
+[32] curated a dataset, known as CVEFixes, from 5365 Common Vulnerabilities and Exposures (CVE) records from the National Vulnerability Database (NVD).
 From each CVE, they automatically extracted the vulnerable and patched versions of each method in open-source projects, along with extensive meta-data such as the corresponding CWEs, project information, and commit data.
-CVEFixes consists of methods extracted from 5495 vulnerability-fixing commits.
-These methods span multiple programming languages such as C/C++, Java, Python, and JavaScript.
-For our work, we extracted all C/C++ and Java methods from CVEFixes.
-We collected 19,576 C/C++ and 3926 Java methods (both vulnerable and non-vulnerable), covering 131 and 68 different CWEs, respectively.
-While many real-world datasets have been proposed in the literature, we needed a dataset for our study that 1) contains vulnerability metadata such as CVE and CWE IDs, 2) is two-sided, i.e., contains both vulnerable and non-vulnerable code samples, and 3) covers multiple languages such as Java and C/C++.
-Table 2 shows a comparison of existing real-world vulnerability datasets.
-We selected CVEFixes because it is the only dataset that fits our criteria.
-| Dataset | Languages | CVE Metadata | Two-Sided | Multi-Lang |
-| BigVul (Fan et al., 2020) | C/C++ | ✓ | ✗ | ✗ |
-| Reveal (Chakraborty et al., 2021) | C/C++ | ✗ | ✗ | ✗ |
-| DiverseVul (Chen et al., 2023) | C/C++ | ✗ | ✓ | ✗ |
-| DeepVD (Wang et al., 2023) | C/C++ | ✗ | ✗ | ✗ |
-| CVEFixes (Bhandari et al., 2021) | C/C++, Java, … | ✓ | ✓ | ✓ |
-2.2. Metrics
+These methods span multiple programming languages but we only consider the C/C++ and Java methods in this work.
+II-B Metrics
 To evaluate the effectiveness of each tool, we use the standard metrics used for classification problems.
@@
 Recall: Recall represents what proportion of vulnerabilities the tool can detect. It is computed as: .
-F1 score: The F1 score is a harmonic mean of precision and recall. It is computed as:
-2.3. Large Language Models
+F1 score: The F1 score is a harmonic mean of precision and recall. It is computed as: .
+II-C Large Language Models
 We choose the most popular state-of-the-art pre-trained Large Language Models (LLMs) for our evaluation.
-We choose OpenAI models: GPT-4 (gpt-4) and GPT-3.5 (gpt-3.5-turbo).
-GPT-3.5 allows up to 4096 input tokens while GPT-4 (which presumably is much larger) allows up to 8192 tokens in the input prompt.
-Since these models are closed-source, we also evaluate CodeLlama models (Rozière et al., 2023), which were recently open-sourced by Meta.
-We select three versions of these models: CodeLlama-7B (CL-7B), CodeLlama-13B (CL-13B), and CodeLlama-34B (CL-34B), containing 7 billion, 13 billion, and 34 billion parameters, respectively.
-We use the Hugging Face APIs (Hugging Face, 2023) to access CodeLlama models.
-We use the “Instruct” version of CodeLlama models—these models are fine-tuned to follow user instructions and hence can better adapt to specific reasoning tasks.
-Table 3 presents LLM details.
-| Model Class | Model Version | Size | Context Window |
+We choose three closed-source models (GPT-4, GPT-3.5 and Gemini-1.5-Flash) and thirteen open-source models from the Codellama-x, Llama-3.1-x, Mistral-Codestral-x, DeepSeekCoder-x, Qwen2.5-x and Qwen2.5-Coder-x series.
+We use the “Instruct” variants of the models wherever applicable since they are fine-tuned to follow user instructions and hence can better adapt to specific reasoning tasks.
+We access the GPT-x models and Gemini-1.5-Flash using the OpenAI and Google Gemini APIs respectively and use the Hugging Face APIs [38] to access the open-source models.
+Table V presents more details about the models.
+| Model | Model Version | Size | Context Size |
+| Qwen-2.5C-1.5B | qwen2.5-coder-1.5b | 1.5B | 128K |
+| Qwen-2.5C-7B | qwen2.5-coder-7b | 7B | 128K |
+| CodeLlama-7B | Codellama-7b-instruct | 7B | 16K |
+| DSCoder-7B | deepseekcoder-7b | 7B | 4K |
+| Llama-3.1-8B | llama-3.1-8b | 8B | 128K |
+| CodeLlama-13B | CodeLlama-13B-Instruct | 13B | 16K |
+| Qwen-2.5-14B | qwen2.5-14b | 14B | 128K |
+| DSCoder-15B | deepseekcoder-v2-15b | 33B | 128K |
+| Codestral-22B | mistral-codestral-22b | 22B | 32K |
+| Qwen-2.5-32B | qwen2.5-32b | 32B | 128K |
+| DSCoder-33B | deepseekcoder-33b | 33B | 16K |
+| CodeLlama-34B | CodeLlama-34B-Instruct | 34B | 16K |
+| Llama-3.1-70B | llama-3.1-70b | 70B | 128K |
+| Gemini-1.5-Flash | gemini-1.5-flash | N/A | 1M |
+| GPT-3.5 | gpt-3.5-turbo-0613 | N/A | 4K |
 | GPT-4 | gpt-4-0613 | N/A | 8K |
-| GPT-3.5 | gpt-3.5-turbo-0613 | N/A | 4k |
-| CodeLlama-34B | CodeLlama-34B-Instruct | 34B | 16k |
-| CodeLlama-13B | CodeLlama-13B-Instruct | 13B | 16k |
-| CodeLlama-7B | CodeLlama-7B-Instruct | 7B | 16k |
-2.4. Prompting Strategies for LLMs
+II-D Prompting Strategies for LLMs
 We explore various prompting strategies that can assist LLMs in predicting if a given code snippet is vulnerable.
-The LLMs discussed in this study support chat interactions with two major types of prompts: the system prompt and the user prompt.
-The system prompt can be used to set the context for the entire conversation while user prompts can be used to provide specific details throughout the chat session.
+The LLMs discussed in this study support chat interactions with two major types of prompts: the system prompt can be used to set the context for the entire conversation while user prompts can be used to provide specific details throughout the chat session.
 We include a system prompt at the start of each input to describe the task and expected structure of the response.
-Since persona assignment has been shown to improve the performance of GPT-4 on specialized tasks (Salewski et al., 2024), we add the line “You are a security researcher, expert in detecting security vulnerabilities” at the start of every system prompt to assign a persona of a Security Researcher to the model.
+Since persona assignment has been shown to improve the performance of GPT-4 on specialized tasks [39], we add the line “You are a security researcher, expert in detecting security vulnerabilities” at the start of every system prompt to assign a persona of a Security Researcher to the model.
 The system prompt for all experiments ends with the statement “Provide response only in the following format:” followed by an expected structure of the response from the model.
@@
 In all our experiments, we incorporate the target code snippet into the user prompt without any changes.
-We construct different prompting strategies to study the reasoning abilities of LLMs:
-2.4.1. Basic prompt
-We design a very simple prompt (shown in Listing 4 in the appendix) with the goal of understanding if the model can take a target code snippet as input and detect if it is vulnerable and determine the correct CWE as well.
+We construct three different prompting strategies:
+II-D1 Basic prompt
+We design a very simple prompt (shown in Listing 4 in the Appendix) to test if the model can take a target code snippet as input and detect if it is vulnerable and determine the correct CWE as well.
 The prompt begins with the message “Is the following code snippet prone to any security vulnerability?” followed by the code snippet.
-| OWASP | Juliet | Juliet | CVEFixes | CVEFixes | Total | |
+| Steps | OWASP | Juliet | Juliet | CVEFixes | CVEFixes | Total |
 | C/C++ | Java | C/C++ | Java | |||
@@
 | Random Selection | 1000 | 1000 | 1000 | 1000 | 1000 | 5000 |
-2.4.2. CWE specific prompt
+II-D2 CWE specific prompt
 The CWE specific prompt is presented in Listing 5 in the Appendix.
 This prompt is similar to the Basic prompt except that it asks the model to predict if the given code snippet is vulnerable to a specific target CWE.
-Hence, the user prompt starts with “Is the following code snippet prone to ¡CWE¿?” followed by the code snippet.
-The <CWE> placeholder here contains both the ID of the CWE and the name.
-For instance, for CWE-22, the user prompt would start with “Is the following code snippet prone to CWE-22 (Improper Limitation of a Pathname to a Restricted Directory (Path Traversal))?” followed by the target code snippet.
-2.4.3. Dataflow analysis-based prompt
-In addition to the straightforward one-step analysis in the previous two prompts, we also study whether providing specific step-by-step analysis instructions can help the LLMs do better reasoning over code and make better predictions.
+Hence, the user prompt starts with “Is the following code snippet prone to CWE?” followed by the code snippet.
+For instance, for CWE-476, the user prompt would start with “Is the following code snippet prone to CWE-476 (NULL Pointer Dereference)?” followed by the target code snippet.
+II-D3 Dataflow analysis-based prompt
 Dataflow analysis is used by several static analysis tools to infer if there exists an unsanitized path from a source to a target node.
-Further, prior literature has shown step-by-step instructions can often elicit better reasoning from LLMs (Wei et al., 2022b).
+Further, prior literature has shown step-by-step instructions can often elicit better reasoning from LLMs [40].
 Motivated by these observations, we designed the CWE-DF prompt (shown in Listing 6 in Appendix) that prompts the model to simulate a source-sink-sanitizer-based dataflow analysis on the target code snippet before predicting if it is vulnerable.
-Naturally, compared to the other prompts, this prompt generates many more tokens and is hence more costly.
-We provide the full prompts in Appendix A.2.
-2.4.4. Other prompting strategies
+Naturally, this prompt is costlier since it generates more tokens than the other prompts.
+We provide the full prompts in Appendix B.
+II-D4 Other prompting strategies
 We also tried other prompting strategies such as Few-shot prompting and Chain-of-thought prompting.
 In the few-shot prompting setup, we include two examples of the task (one with a vulnerability and one without) in the CWE specific prompt before providing the target code snippet.
-Few-shot prompting reported poorer results than the base CWE specific prompt while requiring more tokens.
-Our analysis of the few-shot prompts suggests that providing more examples might not be a useful strategy for vulnerability detection.
-It might be more useful to use prompts that instead elicit reasoning or explanations of some kind before detecting if the given snippet is vulnerable.
-With Chain-of-thought prompting, we explicitly ask the model to provide a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt.
-This setup did not yield better results than the Dataflow analysis-based prompt.
-Moreover, the reasoning chains obtained by Chain-of-thought prompting (both zero-shot and few-shot) were not as elaborate as those from the Dataflow analysis-based prompt thus limiting the ease of debugging.
-We provide the details of this experiment in Appendix A.3.
-2.5. Dataset Processing and Selection
-We perform a data processing and cleaning step for each dataset before evaluating them with LLMs.
-We remove or anonymize information such as commits, benchmark IDs, or vulnerability names that may provide obvious hints about the vulnerability.
-We skip benchmarks that are spread across multiple files, due to limitations of prompt size.
-Table 4 presents the details of our selection stages.
-Due to the prohibitively large cost of running pre-trained LLMs, we select a subset of samples from the original datasets.
-We select samples corresponding to vulnerability types (or CWEs) listed in MITRE’s Top 25 Most Dangerous Software Weaknesses (MITRE Top 25 CWEs, 2023).
-Due to prompt size limitations, we filter out code snippets larger than 2048 tokens.
-Finally, because the datasets are quite large, we further randomly select 500 vulnerable and 500 non-vulnerable samples per dataset.
-Finally, we end up with 5000 samples across five datasets.
-We provide more details for each dataset in Appendix A.1.
-2.6. Experimental Setup
-Experiments with GPT-3.5 and GPT-4.
-We use the OpenAI public API to perform the experiments with GPT-3.5 and GPT-4.
-We use the ChatCompletions API endpoint to query the models with the prompts discussed in Section 2.4.
-We set the sampling temperature to 0 for obtaining deterministic predictions, the maximum number of tokens to 1024, and use the default values for all other parameters.
-In all our experiments, we use the top-1 prediction from the models.
-Experiments with CodeLlama.
-We run all CodeLlama experiments on two sets of machines: one with Intel Xeon machine, with 40 CPUs, four GeForce RTX 2080 Ti GPUs, and 750GB RAM, and another Intel Xeon machine with 48 CPUs, 8 A100s, and 1.5T RAM.
-Similar to GPT-4, we set the temperature to 0, the maximum tokens to 1024, and use top-1 prediction for evaluation.
-3. Results
-3.1. RQ1: Effectiveness of LLMs
-We evaluate the performance of pre-trained LLMs on five open-source datasets discussed in Section 2.1.
-presents the best accuracy and F1 scores (across prompts) of GPT-4, GPT-3.5, and CodeLlama models (CL-7B, CL-13B, CL-34B) on all datasets.
-The more detailed metrics for all prompts are presented in Appendix A.4.
-Modest Vulnerability Detection Performance Across LLMs.
-The average of the highest accuracies across all datasets is 60.6% with the maximum being 77% for CL-7B (with Basic prompt) on the Juliet Java dataset.
-In terms of F1 score, GPT-4 (with CWE-DF) performs the best on the synthetic datasets: OWASP (F1: 0.69), Juliet Java (F1: 0.76), and Juliet C/C++ (F1: 0.70).
-We observe that CodeLlama models, despite being much smaller than GPT-4, are competitive on most datasets, even with the Basic prompt.
-CL-34B reports the highest F1 scores (F1: 0.67) on the real-world datasets, CVEFixes Java and CVEFixes C/C++, with the CWE-DF prompt.
-Performance does not improve with scale.
-While GPT-4 performs strictly better than GPT-3.5 on 3 of the 5 datasets, the CodeLlama models roughly exhibit similar performance.
-Surprisingly, CL-7B performs better than the larger LLMs on the Juliet Java and Juliet C/C++ datasets even with the Basic prompt.
-Listing LABEL:lst:juliet1724 presents a representative example where only CL-7B correctly predicts that an integer overflow vulnerability (CWE-190) cannot occur in the given context.
-Since the data variable can only contain value 2, adding 1 to it in Line LABEL:lst:line:intoverflow:add cannot cause an integer overflow.
-CL-7B recognizes this and predicts that the snippet is not vulnerable.
-CL-13B and GPT-4 (CWE-DF) ignore the hard-coded values and predict that the snippet is vulnerable, indicating that smaller LLMs may sometimes perform better than larger models.
+With Chain-of-thought prompting, we prompt the model to generate a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt.
+Our initial experiments with GPT-4 prompted using these techniques did not yield results better than the Dataflow analysis-based prompt on 100 random samples from two datasets.
+Hence, we do not include these prompts in this study.
+We refer readers to Appendix C for more details.
+II-E Dataset Processing and Selection
+We preprocess each dataset before evaluation by removing or anonymizing information such as commits, benchmark IDs, or vulnerability names that may provide obvious hints about the vulnerability.
+We also skip benchmarks that are spread across multiple files, due to limitations of prompt size.
+We only consider samples corresponding to vulnerability types (CWEs) listed in MITRE’s Top 25 Most Dangerous Software Weaknesses [41].
+We then filter code snippets with more than 2048 tokens due to prompt size limitations and randomly sample 500 vulnerable and 500 non-vulnerable samples per dataset.
+Table VI presents the details of our selection stages.
+We provide more details for each dataset in Appendix A.
+II-F Experimental Setup
+Experiments with closed-source models.
+We use the OpenAI public API’s ChatCompletions endpoint to perform the experiments with GPT-3.5 and GPT-4.
+We use Google’s Gemini API for the experiments with Gemini-1.5-Flash.
+Experiments with open-source models. We run all experiments with the open-source LLMs using the HuggingFace API on a cluster with A100, A6000, and RTX 2080 GPUs.
+In all our experiments, we set the sampling temperature to 0 for obtaining deterministic predictions, the maximum number of tokens to 1024, and use the default values for all other parameters.
+We use the top-1 predictions for evaluation.
+III Results
+III-A RQ1: Effectiveness of LLMs
+We evaluate the performance of pre-trained LLMs on five open-source datasets discussed in Section II-A.
+Figure 1 presents the best accuracy and F1 scores (across prompts) of the 16 models from Table V on all datasets.
+The more detailed metrics for all prompts are presented in Appendix D.
+Finding 1.1: Modest Vulnerability Detection Performance Across LLMs.
+The best performing models and prompts per dataset report an accuracy of 62.8% on average.
+The highest of these is reported by Llama-3.1-70B (with CWE) on the Juliet Java dataset (76%).
+The other best performing models per dataset are: CodeLlama-13B on OWASP (60%), Gemini-1.5-Flash on CVEFixes Java (57%), Qwen-2.5-14B on Juliet C/C++ (65%) and Qwen-2.5-32B on CVEFixes C/C++ (56%).
+This confirms that no model individually performs the best across multiple datasets.
+Moreover, the best accuracies on synthetic datasets are 10.5% higher on average than those on the real-world datasets, suggesting that these models might not be suitable for real-world vulnerability detection yet.
+Finding 1.2: Performance does not improve with scale.
+On the real-world datasets, Qwen-2.5-14B and Qwen-2.5-32B report higher accuracies than the GPT-x models despite being much smaller.
+We see many similar patterns across all models studied implying that model scale does not impact performance.
+Within the same model class, the GPT-x models and the Llama-3.1-x models exhibit improvements in accuracy as the size of the model increases on at least three datasets.
+However, this is not observed in other model classes, i.e., the Qwen-2.5C-x, Qwen-2.5-x, Codellama-x and the DSCoder-x series.
+This is in stark contrast to other domains where increasing model size leads to better performance.
+Listing LABEL:lst:juliet1724 presents a representative example where CodeLlama-7B correctly predicts that an integer overflow vulnerability (CWE-190) cannot occur in the given context while CodeLlama-13B does not.
 [fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]javacode/int_overflow_juliet_java_c13wrong.txt
-3.2. RQ2: Comparison of Prompting Strategies
+III-B RQ2: Comparison of Prompting Strategies
 Figure 2 presents the accuracy and F1 scores (averaged across all LLMs) of the three prompting strategies across all datasets.
-Overall, the three prompts perform similarly in terms of accuracy.
-CWE-DF performs better than CWE and Basic prompt in terms of F1 score on the real-world datasets (CVEFixes C/C++ and CVEFixes Java).
-We next highlight qualitative differences between various prompts:
-The Basic prompt often detects incorrect vulnerability classes.
-Table 5 presents the percentage of samples predicted as vulnerable with the Basic prompt where the predicted CWE is correct, averaged over all datasets with the same programming language.
-We can observe that all models predict an incorrect vulnerability in ¿ 60% and ¿ 53% of all Java and C/C++ samples predicted as vulnerable respectively.
-This suggests that the Basic prompt detects incorrect CWEs in roughly half of the cases.
-| Language (Avg.) | GPT-4 | GPT-3.5 | CL-34B | CL-13B | CL-7B |
-| Java | 0.41 | 0.34 | 0.37 | 0.38 | 0.39 |
-| C/C++ | 0.29 | 0.31 | 0.33 | 0.35 | 0.47 |
-Specifying the CWE in the prompt reduces false alarms.
-In 16 out of the 25 model-dataset combinations, the CWE specific prompt improves or retains both the accuracy and F1 score over the Basic prompt.
-GPT-4 with the CWE specific prompt on Juliet Java reports a 13% higher accuracy and a 0.1 higher F1 score than those with the Basic prompt.
-We manually inspect 10 vulnerable and 10 non-vulnerable samples from this dataset where the CWE specific prompt is correct and the Basic prompt is not.
-We find that not including the CWE in the Basic prompt results in predictions discussing other incorrect CWEs.
-The Basic prompt incorrectly predicts that the 10 non-vulnerable samples are vulnerable to other CWEs of which we only find 3 to be plausible but unlikely due to the context.
+Overall, the three prompts perform similarly in terms of accuracy on all datasets with CWE-DF reporting slightly higher accuracies on the real-world datasets.
+Interestingly, the CWE-DF prompt reports significantly higher F1 scores on average than CWE and Basic prompt on the real-world datasets (by upto 0.18 on CVEFixes C/C++ and 0.14 on CVEFixes Java).
+Furthermore, CWE-DF reports the lowest variance in accuracies and F1 scores across models, as suggested by the standard deviation bars in Figure 2.
+On the other hand, the Basic prompt reports the highest variance in F1 scores.
+We next highlight qualitative differences between prompts:
+Finding 2.1: Specifying the CWE in the prompt reduces false alarms.
+Table VII presents the percentage of samples where the Basic prompt predicts the correct CWE, averaged across datasets for 4 models.
+We observe that the Basic prompt detects incorrect CWEs in 60% and 53% of all Java and C/C++ samples across models.
+We further manually inspected 10 vulnerable and 10 non-vulnerable samples from Juliet Java where GPT-4 with Basic prompt is incorrect.
+In only 4 or the 20 samples, the Basic prompt predicts a plausible CWE.
+However, these CWEs are unlikely due to the context.
 For example, it predicts that a value read from an input stream can be vulnerable if not validated (CWE-20) but this value is not used in a vulnerable context.
-Of the 10 vulnerable samples, Basic prompt predicts a plausible CWE for one sample while incorrectly predicting that the remaining 9 samples are not vulnerable.
-On the other hand, the CWE specific prompt also provides useful high-level explanations for why the snippet is vulnerable/not vulnerable in 18 / 20 samples.
+The CWE specific prompt (i.e., the Basic prompt with CWE) improves or retains accuracy over the Basic prompt on all 5 datasets.
+GPT-4 with the CWE specific prompt not only correctly predicts all the 20 samples discussed above but also provides useful high-level explanations for why the snippet is vulnerable/not vulnerable in 18 / 20 samples.
 The 2 incorrect explanations are artifacts of faulty reasoning or hallucination: for example, an Integer Overflow due to addition to INT_MAX in the function is incorrectly attributed to subtracting from INT_MIN in the explanation.
-Dataflow analysis identifies CWE-relevant textual cues and provides more elaborate explanations.
-The Dataflow analysis-based prompt (CWE-DF) prompt performs better than CWE specific prompt both in terms of accuracy and F1 score in 12 out of 25 model-dataset combinations and reports a higher F1 score in 23 out of 25 cases.
-The most significant improvement with the CWE-DF prompt over the CWE specific prompt is for the CVEFixes Java dataset and GPT-3.5 with 7% higher accuracy and 0.4 higher F1 score.
-We manually inspect 10 vulnerable and 10 non-vulnerable samples from this dataset where only the CWE-DF prompt is correct.
-We find that the CWE-DF prompt leverages textual cues for sanitization (e.g., csrftokenhandler() suggests protection from CSRF) while the CWE specific prompt does not.
-The CWE-DF prompt also assumes that any external parameter or call source is tainted by design and predicts that the sample is vulnerable more often (as is evident from higher F1).
-Further, the analysis from CWE-DF prompt is more detailed and useful in localization than the explanations from CWE specific prompt.
+Based on these observations, specifying the CWE in the prompt can be helpful in reducing incorrect predictions.
+| Language (Avg.) | GPT-4 | GPT-3.5 | CodeLlama-34B | CodeLlama-13B | |
+| Java | 0.41 | 0.34 | 0.37 | 0.38 | |
+| C/C++ | 0.29 | 0.31 | 0.33 | 0.35 |
+Finding 2.2: Dataflow analysis identifies CWE-relevant textual cues and provides actionable explanations.
+The Dataflow analysis-based prompt (CWE-DF) performs better than CWE specific prompt on the real-world datasets, CVEFixes C/C++ and CVEFixes Java.
+We inspect 10 vulnerable and 10 non-vulnerable samples from CVEFixes Java where GPT-3.5 correctly predicts only with CWE-DF.
+We find that the CWE-DF prompt leverages textual cues for sanitization (e.g., csrftokenhandler() suggests protection from CSRF) in 16/20 samples that are missed by CWE specific prompt.
+Further, CWE-DF responses are more useful in localizing the vulnerability as it correctly predicts the correct sources and sinks in 18 / 20 samples, sanitizers in 16 / 20 samples, and unsanitized flows in all samples.
+Listing LABEL:fig:motivating-example presents a response from GPT-4 using CWE-DF prompt that correctly identifies the unsanitized flows between sources and sinks.
+We present more CWE-DF examples in Appendix F.
+Finding 2.3: CWE-DF identifies the correct sources and sinks even when the final prediction is incorrect.
+We also inspect 10 vulnerable and 10 non-vulnerable samples from Juliet C/C++ where CWE-DF’s predictions are incorrect.
+Surprisingly, the sources, sinks and sanitizers are correctly identified in 17 / 20 samples but the unsanitized flows are incorrect in 12 / 20 samples.
+Hence, the final predictions are incorrect only due to erroneous reasoning about the snippet / false assumptions about the CWE.
+Listing LABEL:fig:julietcpp43583 presents an example where the vulnerability is not detected but the sources and sinks are correctly identified.
+This suggests that the CWE-DF prompt can be used to accurately identify sources/sinks/sanitizers and other dataflow analysis techniques can be used to reason about unsanitized flows to predict the vulnerability.
 [fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_51200_highlighted.txt
-3.3. RQ3: Performance of LLMs across CWEs
+III-C RQ3: Performance of LLMs across CWEs
 We next evaluate how the LLMs perform on different classes of security vulnerabilities (CWEs).
+Because the CWE-wise distribution of vulnerable and non-vulnerable samples can be imbalanced, we compute balanced accuracy for each CWE (for ease of presentation, we refer to it as accuracy henceforth in this section).
 For each dataset and model, we consider the best-performing prompt for the analysis and only report CWEs with at least 10 samples.
-Because the CWE-wise distribution of vulnerable and non-vulnerable samples can be imbalanced, we compute balanced accuracy for each CWE (for ease of presentation, we refer to it as accuracy henceforth in this section).
-Figure LABEL:fig:java_cwe_distribution presents the CWE-wise distribution of accuracies on the OWASP, Juliet Java, and CVEFixes Java datasets.
-Figure 3 reports the accuracies on the Juliet C/C++ and CVEFixes C/C++ datasets respectively.
-For each model, we consider the prompt with the best accuracy on the dataset and break ties using the F1 score.
-LLMs perform well on vulnerabilities that do not require additional context.
-We find that there are certain CWEs where multiple LLMs perform better on: Out-of-bounds Read / Write (CWE-125, CWE-787), Null pointer dereference (CWE-476) and Integer Overflow (CWE-190).
+Figure LABEL:fig:java_cwe_distribution presents the CWE-wise distribution of accuracies on the real-world datasets, CVEFixes Java and CVEFixes C/C++.
+Figure 3 reports the accuracies on the synthetic datasets, OWASP, Juliet Java and Juliet C/C++.
+Finding 3.1: LLMs perform well on vulnerabilities that do not require additional context.
+These CWEs include OS Command Injection (CWE-78), Out-of-bounds Read / Write (CWE-125, CWE-787), Null Pointer Dereference (CWE-476), Cross-site Scripting (CWE-79), SQL Injection (CWE-89) and Incorrect Authorization (CWE-863).
 The higher performance on these vulnerabilities can be attributed to the fact that these are fairly self-contained and little additional context is needed to detect them.
-Concretely, GPT-3.5 consistently performs well on NULL Pointer Dereference across all datasets with accuracies of 80.8%, 87.9% and 60.5% on the Juliet C/C++, Juliet Java and CVEFixes C/C++ datasets respectively.
-GPT-4 also performs consistently better on NULL Pointer Dereference and Integer Overflow on the synthetic datasets with accuracies of 62.1% and 68.4% on Juliet C/C++ and 78.5% and 73% on Juliet Java respectively.
-CL-7B reports high accuracies on Integer Overflow on the synthetic datasets with accuracies of 81% on Juliet Java and 70% on Juliet C/C++ respectively.
-GPT-4 and CL-34B perform extremely well on Out-of-bounds Read (78.6% and 64.3%) and Out-of-bounds Write (78.1% and 71.9%) on the CVEFixes Java dataset.
-Poor performance on real-world C/C++ is due to missing global context.
+For example, 4/16 LLMs report accuracies 60% on Incorrect Authorization (CWE-863) on CVEFixes Java it is easier to validate if an implemented authorization check is correct or not.
+On the other hand, no LLMs report high accuracies on Missing Authorization (CWE-862) since it’s not known if an input parameter has been authorized outside the context of the target method and more context is hence needed to detect this vulnerability class.
+The following summarizes how LLMs perform on these CWEs:
+OS Command Injection (CWE-78) sees the highest performance across models and datasets with 60% accuracies reported by 5/16 LLMs on CVEFixes Java and 11/16 LLMs on CVEFixes C/C++.
+Llama-3.1-70B reports the best performance on CWE-78 with accuracies 64% and 78% on CVEFixes Java and CVEFixes C/C++ respectively.
+Llama-3.1-8B and GPT-4 perform extremely well on Out-of-bounds Write (CWE-787) with accuracies of 89% and 79% on CVEFixes Java and CVEFixes C/C++ respectively and on Out-of-bounds Read (CWE-125) with accuracies of 84% and 78% respectively.
+Moreover, accuracies over 60% are reported on CVEFixes Java by 5/16 LLMs for CWE-125 and 4/16 LLMs for CWE-787.
+NULL Pointer Dereference (CWE-476) sees accuracies higher than 60% by 3/16 LLMs on CVEFixes C/C++, 7/16 on Juliet C/C++ and 11/16 on OWASP.
+Interestingly, DSCoder-7B performs the best on all three datasets with accuracies of 63%, 88% and 83% respectively.
+GPT-3.5 reports the highest accuracy of 70% on SQL Injection (CWE-89) on CVEFixes Java and 7/16 LLMs report accuracies over 60%.
+Surprisingly, all LLMs report accuracies 60% on the same CWE on synthetic datasets (OWASP and Juliet Java).
+Qwen-2.5-32B reports high accuracies of 67% and 66% on Cross-Site Scripting (CWE-79) on OWASP and CVEFixes Java respectively.
+Accuracies over 60% are reported by 3/16 LLMs on CVEFixes Java and 11/16 LLMs on OWASP.
+Finding 3.2: Poor performance on real-world C/C++ is due to missing global context.
 We see that the performance of all LLMs on vulnerabilities in CVEFixes C/C++ is worse than that on the same CWEs in CVEFixes Java and Juliet C/C++.
-For instance, while GPT-4 and CL-34B perform extremely well on the Out-of-bounds Read / Write vulnerabilities in CVEFixes Java as discussed above, they report accuracies ¡ 53% for these CWEs on the CVEFixes C/C++ dataset.
+While GPT-4 and Llama-3.1-8B perform extremely well on the Out-of-bounds Read / Write vulnerabilities in CVEFixes Java as discussed above, they report accuracies 53% for these CWEs on the CVEFixes C/C++ dataset.
+In fact, no LLMs report accuracies 60% for these CWEs on CVEFixes C/C++.
 We attribute this disparity to the nature of these vulnerabilities in the two languages: Out-of-bounds Reads / Writes in CVEFixes C/C++ require reasoning about pointers and structs which requires more context about the structs and their members.
@@
 This issue does not emerge in Juliet C/C++ because all the information about the pointers is presented in the snippet.
-We present examples in Appendix A.8.
-Some LLMs are better at detecting certain CWEs.
-GPT-4 reports high accuracies on CWE-78 (OS Command Injection) in 3/5 datasets (CVEFixes C/C++ with 70.8%, CVEFixes Java with 61.7%, and Juliet Java with 63.8%) but the other models do not.
-Among the CodeLlama models, CL-13B performs well on CWE-362 (Race Condition) in CVEFixes C/C++ with 63.1% and CVEFixes Java with 66.7%.
-CWE-89 (SQL Injection) is another vulnerability where only CL-7B reports high accuracies on 2/3 datasets (Juliet Java with 74% and CVEFixes Java with 64%).
-3.4. RQ4: Impact of Adversarial Attacks
+We present examples in Appendix G.
+Finding 3.3: Some LLMs are better at detecting certain CWEs. Concretely, the following LLMs report the best accuracies on certain CWEs across datasets:
+Llama-3.1-70B on OS Command Injection (CWE-78)
+DSCoder-7B on NULL Pointer Dereference (CWE-476)
+Qwen-2.5-32B on Cross-Site Scripting (CWE-79)
+Llama-3.1-8B and GPT-4 on Java Out-of-bounds Read/Write (CWE-125/787)
+III-D RQ4: LLMs vs Static Analysis Tools
 Experiment setup.
-We implement three semantics-preserving attacks, proposed by Gao et al.
-(Gao et al., 2023): dead-code injection, variable renaming, and branch insertion (illustrated in Table 11 in appendix).
-We only select test samples where the original prediction of LLMs (with the CWE prompt) was correct.
-We select 100 test samples per dataset for this experiment and measure the reduction in accuracy after the attack.
-For this experiment, we select two models, GPT-4 and CL-7B, and four datasets.
-Figure 4 presents our results.
-Results.
-We find GPT-4 suffers mild degradation in accuracy (average 8.6%) across all datasets and attacks.
-For GPT-4, the degradation is more significant for real-world datasets (up to 23%) compared to synthetic ones (up to only 2%).
-In contrast, for CodeLlama-7B, the accuracy significantly reduces by almost 40-50% in many cases and by 39% on average, especially for C++.
-Among the three attacks, branch insertion (27% avg.
-reduction) and variable renaming (23% avg.
-reduction) have the highest impact on LLM performance.
-On further analysis of incorrect predictions, we observe that under these attacks, GPT-4 either fails to infer any sources, sinks, or sanitizers (even if they exist) or the reasoning chain breaks (which worked previously).
-Interestingly, under these attacks, CodeLlama-7B fails to follow the instructions in the system prompt and does not perform proper analysis using the dataflow analysis steps, causing it to output an incorrect response.
-3.5. RQ5: LLMs vs Static Analysis Tools
+We next explore how the LLMs compare against CodeQL.
+Since CodeQL requires building projects before analysis and the real-world datasets contain large projects, we limit our focus to the three synthetic datasets, namely OWASP, Juliet Java and Juliet C/C++.
+We use the official CodeQL queries designed for the top 25 CWEs.
+Table VIII presents results from CodeQL and the best performing LLM on the three datasets: CodeLlama-13B on OWASP, Llama-3.1-70B on Juliet Java and Qwen-2.5-14B on Juliet C/C++.
+Finding 4.1: CodeQL performs better than the LLMs in terms of accuracy on all three datasets.
+CodeQL reports 0.07 and 0.15 higher F1 than CodeLlama-13B on OWASP and Llama-3.1-70B on Juliet Java respectively.
+However, Qwen-2.5-14B reports a 0.11 higher F1 on Juliet C/C++.
+Finding 4.2: LLMs perform better than CodeQL on certain CWEs.
+A CWE-wise comparison of LLMs with CodeQL in Figure 3 reveals that LLMs report higher accuracies on CWE-22 (Path Traversal), CWE-78 (OS Command Injection), CWE-476 (NULL Pointer Dereference), 416 (Use After Free) on at least 2 / 3 datasets while CodeQL performs better on CWE-190 (Integer Overflow) on 2 datasets.
+Concretely,
+CodeQL performs better than the LLMs on CWE-190 (Integer Overflow) with 11% and 21% higher accuracies than Llama-3.1-70B on Juliet Java and Juliet C/C++ respectively.
+On the other hand, DSCoder-7B performs better than CodeQL on CWE-476 (NULL Pointer Dereference) with 12% higher accuracy on Juliet Java and only 1% lower accuracy on Juliet C/C++.
+Moreover, 6 / 16 LLMs report accuracies higher than CodeQL on CWE-476 from Juliet Java.
+While CodeQL reports an extremely high accuracy of 95% on CWE-78 (OS Command Injection) with Juliet Java, it is outperformed by CodeLlama-13B by 1% and Codestral-22B by 7% on OWASP and Juliet C/C++ respectively.
+Interestingly, 7 / 16 LLMs report higher accuracies (by upto 10%) at detecting CWE-22 (Path Traversal) on OWASP.
+Similarly, 3 LLMs perform better on CWE-416 (Use After Free) from Juliet C/C++ with DSCoder-7B reporting a whopping 24% higher accuracy than CodeQL.
+Finding 4.3: CodeQL’s worse performance on some CWEs can be attributed to the very specific manually-written queries which may not cover all possible scenarios of the vulnerability.
+For example, CodeQL only detects CWE-78 (OS Command Injection) in C/C++ snippets when there exist system commands that take a string of arguments (execl).
+This query cannot handle commands that take a list of arguments (eg., _execvp).
+Listing LABEL:fig:motivating-example provides an example of this scenario where CodeQL does not detect that the snippet is prone to OS Command Injection but GPT-4 (CWE-DF) correctly identifies _execvp as a vulnerable sink.
+Listing LABEL:fig:julietcpp43583 presents an example where CodeQL correctly predicts that the target snippet is vulnerable to Integer Overflow while GPT-4 with CWE-DF does not.
+However, the model correctly identifies the sources, sinks and even unsanitized dataflows in this case but fails to faithfully reason over them when predicting the vulnerability.
+These examples support the observation from Section III-B that LLMs can be used to identify sources and sinks relevant to the target vulnerability (which can be missed by manually written queries) while static analysis tools like CodeQL can be used to reason over them.
+More examples are presented in the Appendix E.
+| Dataset/Model | CodeQL | Best performing LLM | ||||||
+| A | P | R | F1 | A | P | R | F1 | |
+| OWASP | 0.63 | 0.58 | 0.96 | 0.72 | 0.60 | 0.58 | 0.74 | 0.65 |
+| Juliet Java | 0.94 | 0.92 | 0.95 | 0.94 | 0.76 | 0.71 | 0.89 | 0.79 |
+| Juliet C/C++ | 0.72 | 0.99 | 0.44 | 0.61 | 0.65 | 0.60 | 0.89 | 0.72 |
+[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_43583_highlighted.txt
+III-E RQ5: LLMs vs Deep Learning-Based Tools
+We compare LLMs against two prior deep learning-based approaches: 1) DeepDFA [10], which trains Graph Neural Networks using embeddings of control flow graphs and associated data flow facts, and 2) LineVul [11], which is a transformed-based model trained using token-based representation of code.
 Experiment setup.
-We next explore how GPT-4 (CWE-DF) compares against CodeQL.
-Since CodeQL requires building projects before analysis and the real-world datasets contain large projects, we limit our focus to the three synthetic datasets, namely OWASP and Juliet Java and Juliet C/C++.
-In this study, we run the official CodeQL queries designed for the top 25 CWEs.
-Table 6 presents results from CodeQL and GPT-4 on the three datasets.
-CodeQL performs better than GPT-4 (CWE-DF) in terms of accuracy on all three datasets.
-GPT-4 reports a 0.09 higher F1 than CodeQL on Juliet C/C++ while CodeQL reports a 0.03 and 0.18 higher F1 on OWASP and Juliet Java respectively.
-| Dataset/Model | CodeQL | GPT-4 (CWE-DF) | ||||||
-| A | P | R | F1 | A | P | R | F1 | |
-| OWASP | 0.63 | 0.58 | 0.96 | 0.72 | 0.55 | 0.53 | 1.00 | 0.69 |
-| Juliet Java | 0.94 | 0.92 | 0.95 | 0.94 | 0.70 | 0.63 | 0.98 | 0.76 |
-| Juliet C/C++ | 0.72 | 0.99 | 0.44 | 0.61 | 0.59 | 0.55 | 0.98 | 0.70 |
-[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_43583_highlighted.txt
-3.6. RQ6: LLMs vs Deep-Learning-Based Tools
-We compare LLMs against two prior deep learning-based approaches: 1) DeepDFA (Steenhoek et al., 2024), which trains Graph Neural Networks using embeddings of control flow graphs and associated data flow facts, and 2) LineVul (Fu and Tantithamthavorn, 2022), which is a transformed-based model trained using token-based representation of code.
-Experiment setup.
-We used the DeepDFA and LineVul versions from DeepDFA’s latest artifact version (DeepDFA artifact, 2024).
-For CVEFixes C/C++ and Juliet C/C++, we use the same test set as our main evaluation §3.1.
-We split the remaining dataset into training (80%) and validation (20%) sets.
-We also compare the results with the BigVul (Fan et al., 2020) dataset, used in the prior work.
-Because BigVul is a much larger real-world dataset, it can potentially improve the performance of the tools.
-We explore two evaluation settings: 1) train and test on the same dataset and 2) train on one and test on a different dataset to evaluate the generalizability of these models.
-Table 7 presents the results.
-Each cell reports the average results across three runs.
-DL-based approaches have limited effectiveness on real-world datasets.
-We observe that both DeepDFA and LineVul obtain a maximum F1 score of 0.62 on CVEFixes C/C++, which is quite similar to the scores of GPT-4.
-For DeepDFA, while training on larger BigVul and Juliet C/C++ datasets improves its F1 scores, the accuracy scores remain close to 50%.
-On the other hand, LineVul is able to learn better even when using CVEFixes C/C++, which is relatively smaller.
-Interestingly, when trained on BigVul, LineVul obtains a poor F1 score of 0.02 on CVEFixes C/C++, which might indicate an over-fitting problem.
-We observe a similar pattern when LineVul is tested on Juliet C/C++ using CVEFixes C/C++ or BigVul training datasets.
-For Juliet C/C++, both tools obtain better scores compared to GPT-4.
-LineVul obtains a perfect score indicating that Juliet C/C++ has limited code patterns that are likely easier to learn.
-However, as we explain later, LineVul does not generalize well.
-Generalizability across datasets.
-We observe that, for LineVul, the F1 scores drop drastically when trained on BigVul and tested on CVEFixes C/C++ or Juliet C/C++.
-Compared to LineVul, DeepDFA generalizes better, obtaining higher accuracy and F1 scores on Juliet C/C++ when trained using other datasets.
-Trade-offs.
+Our main aim in this experiment is to check the generalizability of these techniques on real-world datasets beyond the datasets they are trained on.
+We use CVEFixes C/C++ with the 1000 samples from our main evaluation §III-A as the real-world test set.
+We train on three different datasets:Juliet C/C++, CVEFixes C/C++ (excluding samples in the test set) and BigVul [34](the original C/C++ dataset that these models were trained on).
+We use an 80/20 train-validation split while training on these datasets.
+We use the DeepDFA and LineVul versions from DeepDFA’s latest artifact [42].
+Table IX presents the results, averaged across three runs.
+Finding 5.1: DL-based approaches have limited effectiveness on real-world datasets.
+DeepDFA and LineVul trained on CVEFixes C/C++ training set report accuracies of 51% and 59% on CVEFixes C/C++ test set respectively while the best performing LLM on this dataset, Qwen-2.5-32B, reports an accuracy of 56%.
+This is in stark contrast to the performance of these techniques on BigVul where they report accuracies higher than 98%.
+Moreover, Qwen-2.5-32B reports an F1 score of 0.65 which is 0.42 and 0.04 higher than DeepDFA and LineVul respectively.
+Finding 5.2: DL-based approaches do not generalize across datasets.
+When trained on Juliet C/C++ or BigVul, both DeepDFA and LineVul report accuracies and F1 scores lower than Qwen-2.5-32B by upto 6% and 0.63 respectively.
+Finding 5.3: LLMs are preferable over DL-based approaches due to low inference overheads and natural language explanations.
 DeepDFA involves significant inference overhead, due to the CFG extraction and dataflow analysis steps.
@@
 | LineVul | BigVul | BigVul | 0.99 | 0.96 | 0.88 | 0.92 |
-| GPT-4 | CWE-DF | CVEFixes C/C++ | 0.52 | 0.51 | 0.76 | 0.61 |
+| Qwen-2.5-32B | CWE-DF | CVEFixes C/C++ | 0.56 | 0.54 | 0.81 | 0.65 |
 | DeepDFA | CVEFixes C/C++ | CVEFixes C/C++ | 0.51 | 0.55 | 0.17 | 0.23 |
@@
 | LineVul | BigVul | CVEFixes C/C++ | 0.50 | 0.63 | 0.01 | 0.02 |
-| GPT-4 | CWE-DF | Juliet C/C++ | 0.59 | 0.55 | 0.98 | 0.70 |
-| DeepDFA | Juliet C/C++ | Juliet C/C++ | 0.77 | 0.74 | 0.82 | 0.78 |
-| DeepDFA | CVEFixes C/C++ | Juliet C/C++ | 0.64 | 0.72 | 0.46 | 0.55 |
-| DeepDFA | BigVul | Juliet C/C++ | 0.74 | 0.70 | 0.82 | 0.77 |
-| LineVul | Juliet C/C++ | Juliet C/C++ | 1.0 | 1.0 | 0.99 | 1.0 |
-| LineVul | CVEFixes C/C++ | Juliet C/C++ | 0.51 | 0.62 | 0.14 | 0.22 |
-| LineVul | BigVul | Juliet C/C++ | 0.43 | 0.42 | 0.39 | 0.41 |
-4. Related Work
+IV Related Work
 Static analysis tools for vulnerability detection.
-Static analysis tools search for pre-defined vulnerability patterns in code.
-Tools such as FlawFinder (FlawFinder, 2023) and CppCheck (CPPCheck, 2023) use syntactic and simple semantic analysis techniques to find vulnerabilities in C++ code.
-More advanced tools like CodeQL (Avgustinov et al., 2016), Infer (Fb Infer, 2023), and CodeChecker (Code Checker, 2023) employ semantic analysis techniques and can detect vulnerabilities in multiple languages.
+Tools such as FlawFinder [43] and CppCheck [44] use syntactic and simple semantic analysis techniques to find vulnerabilities in C++ code.
+More advanced tools like CodeQL [5], Infer [45], and CodeChecker [46] employ semantic analysis techniques and can detect vulnerabilities in multiple languages.
 Static analysis tools rely on manually crafted rules and precise specifications of code behavior, which is difficult to obtain automatically.
-In contrast, while LLMs cannot always reliably perform end-to-end reasoning over code, we observe that LLMs can automatically identify such specifications by leveraging statistically learned rules from training data.
-Further, due to their semantic understanding of natural language, LLMs can also perform more contextual reasoning.
-These abilities can potentially be leveraged to improve static analysis tools.
+In contrast, while LLMs cannot always reliably perform end-to-end reasoning over code, we find that they are capable of automatically identifying these specifications which can be leveraged to improve static analysis-based detection tools.
 Deep Learning-based vulnerability detection.
 Several works have focused on using Deep Learning techniques for vulnerability detection.
-Earlier works such as Devign (Zhou et al., 2019), Reveal (Chakraborty et al., 2020), LineVD (Hin et al., 2022) and IVDetect (Li et al., 2021) leveraged Graph Neural Networks (GNNs) for modeling dataflow graphs, control flow graphs, abstract syntax trees and program dependency graphs.
-Other works explored alternate model architectures: VulDeePecker (Li et al., 2020) and SySeVR (Li et al., 2018) used LSTM-based models on slices and data dependencies while Draper used Convolutional Neural Networks.
-Recent works demonstrate that Transformer-based models fine-tuned on the task of vulnerability detection can outperform specialized techniques (CodeBERT, LineVul (Fu and Tantithamthavorn, 2022), UnixCoder).
-DeepDFA (Steenhoek et al., 2024) and ContraFlow (Cheng et al., 2022) learn specialized embeddings that can further improve the performance of Transformer-based vulnerability detection tools.
-To the best of our knowledge, these techniques provide binary labels for vulnerability detection and cannot classify the type of vulnerability.
-Thapa et al.
-(2022) explore whether Language Models fine-tuned on multi-class classification can perform well where the classes correspond to groups of similar types of vulnerabilities.
-In contrast, we study some of the largest Language Models, such as GPT-4, and perform a much granular CWE-level classification, generate human-readable informal specifications and explore various prompting techniques that allow using the LLMs out-of-the-box.
+Earlier works such as Devign [47], Reveal [48], LineVD [49] and IVDetect [50] leveraged Graph Neural Networks (GNNs) for modeling dataflow graphs, control flow graphs, abstract syntax trees and program dependency graphs.
+Other works explored alternate model architectures: VulDeePecker [51] and SySeVR [52] used LSTM-based models on slices and data dependencies while Draper used Convolutional Neural Networks.
+Recent works demonstrate that Transformer-based models fine-tuned on the task of vulnerability detection can outperform specialized techniques (CodeBERT, LineVul [11], UnixCoder).
+DeepDFA [10] and ContraFlow [53] learn specialized embeddings that can further improve the performance of Transformer-based vulnerability detection tools.
+These techniques, however, provide binary labels for vulnerability detection and do not provide natural language explanations.
 LLMs for automated software engineering.
-Several recent approaches have demonstrated that LLMs can be effectively leveraged to improve the state-of-the-art performance in various traditional software engineering tasks such as automated program repair (Xia et al., 2023; Joshi et al., 2023; Xia and Zhang, 2022), test generation (Lemieux et al., 2023; Deng et al., 2023), code evolution (Zhang et al., 2023), and fault localization (Yang et al., 2024).
-However, unlike these approaches, we find that LLMs have limited vulnerability detection abilities.
-Recently, Li et al.
-(Li et al., 2024) developed Llift, an approach that combines LLMs with static analysis to detect Use Before Initialization (UBI) bugs in Linux kernel.
-While they focus on a specific class of bugs, their approach supports the observations we make in Section 3.5 on the complementary nature of static analysis and LLMs.
-To the best of our knowledge, our work is the first comprehensive and general study of vulnerability detection abilities of LLMs across a broad range of vulnerabilities (25 CWEs) across two languages.
-5. Threats to Validity
+Recent approaches have demonstrated the effectiveness of LLMs on software engineering tasks such as automated program repair [13, 14, 15], test generation [16, 17], code evolution [18], and fault localization [19].
+However, unlike these approaches, we find that scaling LLMs to larger sizes does not improve vulnerability detection abilities.
+[54] explore whether Language Models fine-tuned on multi-class classification can perform well where the classes correspond to groups of similar types of vulnerabilities.
+In contrast, we study whether LLMs can perform a much granular CWE-level classification through prompting.
+Recent work combining LLMs with static analysis to detect Use Before Initialization (UBI) bugs in the Linux Kernel [55] supports our claims in Section III-D (but focuses on a specific class of bugs).
+There are other concurrent studies evaluating LLMs on vulnerability detection [24, 25, 26, 27, 28].
+Table II provides a comparison against these studies.
+Section III-A corroborates the common finding from these studies that LLMs do not perform well on real-world datasets.
+However, to the best of our knowledge, our study is the first work that qualitatively identifies prompts and vulnerability classes where LLMs perform well (and often better than other tools).
+Moreover, our study focuses on a larger/different class of state-of-the-art LLMs, datasets, languages, and vulnerability classes.
+V Threats to Validity
 External.
@@
 We also choose both state-of-the-art closed-source and open-source LLMs.
-However, our insights may not generalize to other languages or datasets not studied in this paper.
+However, our insights may not generalize to other languages or datasets.
 Internal.
 Owing to the non-deterministic nature of LLMs and single experiment runs per benchmark, our observations may be biased.
-To mitigate this threat, we use a temperature of zero to ensure determinism across all LLMs.
+To mitigate this threat, we use a temperature of 0 to ensure determinism across all LLMs.
 While this works well for locally run CodeLlama models, it is well-known that GPT-4 and GPT-3.5 might still return non-deterministic results.
-However, due to the large number of benchmarks we evaluate, the non-determinism should balance out across the datasets.
-Further, given the poor effectiveness of all LLMs across the board, we do not expect our results to significantly change with re-runs.
+However, this should balance out across datasets and over the large number of benchmarks we evaluate on.
+Further, given similar results across LLMs on real-world-datasets, we do not expect significant changes with re-runs.
 Our evaluation code and scripts may have bugs, which might bias our results.
@@
 To address this threat, multiple co-authors reviewed the code regularly and actively fixed issues.
-Further, multiple co-authors of the paper independently analyzed the results and discussed them together to mitigate any discrepancies.
-6. Conclusion
+Further, multiple co-authors independently analyzed the results and discussed them together to mitigate any discrepancies.
+VI Conclusion
 In this work, we performed a comprehensive analysis of LLMs for security vulnerability detection.
 Our study reveals that both closed-source LLMs, such as GPT-4, and open-source LLMs, like CodeLlama, perform modestly at vulnerability detection for both Java and C/C++.
-Their performance on vulnerability detection is lower when dealing with real-world code.
-However, we find that even in cases where the models produce incorrect predictions, they identify relevant sources, sinks and sanitizers for dataflow analysis.
+However, we find specific vulnerability classes where LLMs excel (often performing better than static analysis tools, such as CodeQL).
+Moreover, we find that even in cases where the models produce incorrect predictions, they are capable of identifying relevant sources, sinks and sanitizers that can be used for downstream dataflow analysis.
 Hence, we believe that an interesting future direction is to develop neuro-symbolic techniques that combine the intuitive reasoning abilities of LLMs with symbolic tools such as logical reasoning engines and static code analyzers for more effective and interpretable solutions.
-Appendix A Appendix
-A.1. Dataset Processing and Selection
-We perform a data processing and cleaning step for each dataset before evaluating them with LLMs.
-OWASP.
-We remove or anonymize information in OWASP benchmarks that may provide obvious hints about the vulnerability in a file.
-For instance, we change package, variable names, and strings such as “owasp”, “testcode”, and “/sqli-06/BenchmarkTest02732” to other pre-selected un-identifying names such as “pcks”, “csdr”, etc.
-We remove all comments in the file because they may explicitly highlight the vulnerable line of code or may have irrelevant text (such as copyright info), which may leak information.
-These changes, however, do not change the semantics of the code snippets.
-Juliet Java and C/C++.
-Similar to OWASP, we remove all comments and transform all identifiers that leak identifying information in all test cases in the Juliet benchmark.
-For instance, we change “class CWE80_XSS_CWE182_Servlet_connect_tcp_01” to “class MyClass”.
-The Juliet benchmark provides the vulnerable (named as “bad”) and non-vulnerable (named as “good*”) methods in the same file.
-For easier evaluation, we perform a pre-processing step to split each file into two, each containing either a vulnerable or non-vulnerable method.
-Juliet also contains special benchmarks that have dependencies across multiple (2-5) files.
-We skip these benchmarks because they are typically too big to fit into the LLM prompt.
-Hence, the number of test cases after the data processing step in Juliet is reduced (as shown in Table 4).
-CVEFixes.
-For each CVE, CVEFixes provides the methods that were involved in the fix commit.
-It also includes the method code in the parent commit, i.e., the method version before the fix.
-We collect all methods in the fix commit and the parent commit and label them as vulnerable and non-vulnerable, respectively.
-Similar to other datasets, we also remove all comments in the method code.
-While CVEFixes contains methods across multiple programming languages, we only collect C/C++ and Java methods for our study.
-A.2. Prompting Strategies
-The Basic prompt is presented in Listing 1, CWE specific prompt in Listing 2 and Dataflow analysis-based prompt in Listing 3.
-A.3. Other Prompting Strategies
-In addition to the prompting strategies presented in our main evaluation, we considered other popular prompting strategies such as Few-shot prompting and Chain-of-thought prompting in a limited experimental setting.
-For the few-shot prompt (CWE-Few-shot), we included two examples of the task (one with a vulnerability and one without) in the CWE specific prompt before providing the target code snippet.
-For the chain-of-thought prompt (CWE-CoT), we explicitly ask the model to provide a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt.
-Table 8 and Table 9 present the results from GPT-4 with various prompting strategies on a random subset of 100 samples of the Juliet Java and CVEFixes C/C++ datasets respectively.
-The CWE-DF prompt reports the highest accuracy of 69% and the highest F1 score of 0.75 on the Juliet Java dataset.
-The CWE-DF prompt reports a 0.05 higher F1 score than the CWE-CoT prompt and a 0.03 higher F1 score than the CWE-Few-shot prompt.
-This difference is much more prominent on the CVEFixes C/C++ dataset where the CWE-DF prompt reports a 0.34 higher F1 score than the CWE-CoT prompt and a 0.31 higher F1 score than the CWE-Few-shot prompt.
-Moreover, the CWE-Few-shot prompt reported a 0.2 lower F1 score than the CWE specific prompt on the CVEFixes C/C++ dataset while requiring more tokens.
-Our analysis of the few-shot prompts suggests that providing more examples may not be a useful strategy for vulnerability detection.
-Because the potential set of vulnerable code patterns is quite large, the provided examples hardly make a difference to LLMs’ reasoning abilities.
-Hence, it may be more useful to use prompts that instead elicit reasoning or explanations of some kind before detecting if the given snippet is vulnerable.
-The CWE-CoT prompt, however, does not help with reasoning always, as it either performed at par or worse than the Dataflow analysis-based prompt.
-| Model | Prompt | Metrics | ||||
-| A | P | R | F1 | |||
-| GPT-4 | CWE | 0.65 | 0.58 | 0.96 | 0.72 | |
-| GPT-4 | CWE-Few-shot | 0.65 | 0.58 | 0.94 | 0.72 | |
-| GPT-4 | CWE-CoT | 0.69 | 0.64 | 0.79 | 0.70 | |
-| GPT-4 | CWE-DF | 0.69 | 0.61 | 0.96 | 0.75 | |
-| Model | Prompt | Metrics | |||||
-| A | P | R | F1 | ||||
-| GPT-4 | CWE | 0.55 | 0.54 | 0.58 | 0.56 | ||
-| GPT-4 | CWE-Few-shot | 0.49 | 0.38 | 0.34 | 0.36 | ||
-| GPT-4 | CWE-CoT | 0.52 | 0.37 | 0.30 | 0.33 | ||
-| GPT-4 | CWE-DF | 0.56 | 0.56 | 0.83 | 0.67 | ||
-Learning from these experiments, we selected the CWE specific prompt, Dataflow analysis-based prompt, in addition to the Basic prompt, for our main evaluation with LLMs.
-A.4. Detailed metrics across all LLMs and Datasets
-Table 10 presents the metrics for all LLMs and Datasets across all prompts.
-| Model | Prompt | OWASP | Juliet Java | CVEFixes Java | Juliet C/C++ | CVEFixes C/C++ | ||||||||||||||||||||
-| A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | |||||||
-| GPT-4 | Basic | 0.52 | 0.51 | 1.00 | 0.68 | 0.56 | 0.54 | 0.85 | 0.66 | 0.50 | 0.50 | 0.34 | 0.41 | 0.54 | 0.52 | 0.92 | 0.67 | 0.51 | 0.50 | 0.57 | 0.54 | |||||
-| GPT-4 | CWE | 0.54 | 0.52 | 1.00 | 0.69 | 0.69 | 0.63 | 0.97 | 0.76 | 0.55 | 0.56 | 0.44 | 0.50 | 0.58 | 0.54 | 0.95 | 0.69 | 0.52 | 0.52 | 0.52 | 0.52 | |||||
-| GPT-4 | CWE-DF | 0.55 | 0.53 | 1.00 | 0.69 | 0.70 | 0.63 | 0.98 | 0.76 | 0.53 | 0.53 | 0.59 | 0.56 | 0.59 | 0.55 | 0.98 | 0.70 | 0.52 | 0.51 | 0.76 | 0.61 | |||||
-| GPT-3.5 | Basic | 0.53 | 0.52 | 0.72 | 0.60 | 0.58 | 0.57 | 0.71 | 0.63 | 0.46 | 0.35 | 0.09 | 0.15 | 0.49 | 0.49 | 0.64 | 0.56 | 0.52 | 0.55 | 0.19 | 0.29 | |||||
-| GPT-3.5 | CWE | 0.55 | 0.54 | 0.62 | 0.58 | 0.52 | 0.52 | 0.55 | 0.54 | 0.47 | 0.41 | 0.12 | 0.19 | 0.49 | 0.49 | 0.70 | 0.58 | 0.51 | 0.54 | 0.19 | 0.28 | |||||
-| GPT-3.5 | CWE-DF | 0.51 | 0.51 | 0.93 | 0.66 | 0.40 | 0.44 | 0.73 | 0.55 | 0.54 | 0.53 | 0.66 | 0.59 | 0.40 | 0.44 | 0.77 | 0.56 | 0.52 | 0.52 | 0.75 | 0.61 | |||||
-| CL-34B | Basic | 0.51 | 0.51 | 1.00 | 0.67 | 0.47 | 0.48 | 0.85 | 0.62 | 0.50 | 0.50 | 0.28 | 0.36 | 0.50 | 0.50 | 0.93 | 0.65 | 0.51 | 0.51 | 0.19 | 0.28 | |||||
-| CL-34B | CWE | 0.57 | 0.54 | 0.94 | 0.69 | 0.49 | 0.49 | 0.94 | 0.65 | 0.50 | 0.51 | 0.16 | 0.25 | 0.53 | 0.52 | 0.98 | 0.68 | 0.51 | 0.54 | 0.08 | 0.14 | |||||
-| CL-34B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | |||||
-| CL-13B | Basic | 0.52 | 0.51 | 0.98 | 0.67 | 0.47 | 0.48 | 0.72 | 0.58 | 0.50 | 0.50 | 0.13 | 0.21 | 0.46 | 0.48 | 0.79 | 0.59 | 0.51 | 0.52 | 0.22 | 0.31 | |||||
-| CL-13B | CWE | 0.52 | 0.51 | 0.98 | 0.67 | 0.50 | 0.50 | 0.89 | 0.64 | 0.48 | 0.47 | 0.29 | 0.36 | 0.53 | 0.51 | 0.98 | 0.67 | 0.52 | 0.52 | 0.56 | 0.54 | |||||
-| CL-13B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 0.96 | 0.66 | |||||
-| CL-7B | Basic | 0.56 | 0.54 | 0.81 | 0.65 | 0.77 | 0.79 | 0.72 | 0.75 | 0.47 | 0.46 | 0.42 | 0.44 | 0.62 | 0.58 | 0.86 | 0.69 | 0.47 | 0.35 | 0.07 | 0.11 | |||||
-| CL-7B | CWE | 0.50 | 0.50 | 1.00 | 0.67 | 0.52 | 0.51 | 0.99 | 0.67 | 0.51 | 0.51 | 0.84 | 0.63 | 0.51 | 0.50 | 0.99 | 0.67 | 0.50 | 0.50 | 0.84 | 0.63 | |||||
-| CL-7B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.51 | 0.50 | 0.98 | 0.66 | |||||
-A.5. Adversarial Attack Examples
-| Attack | Example |
-| Dead-code Injection | ⬇ int obj = new Object(...); if (obj == null) { /* deadcode */ } |
-| Variable Renaming | ⬇ String var3 = (String) names.nextElement(); String[] values = request.getParameterValues(var3); |
-| Branch Insertion | ⬇ boolean var11 = true; if (var11) { configManager.init(); } |
-Table 11 presents the three adversarial attacks that we implemented: dead-code injection, variable renaming, and branch insertion.
-Each attack contains multiple code patterns, which are randomly chosen when perturbing a given code snippet.
-We implement these attacks for Java and C/C++.
-For a given code snippet, we only apply an attack only once.
-A.6. More Details on CodeQL vs. GPT-4
-Table 12 presents the CWE-wise distribution of CodeQL results.
-CodeQL shows relatively better performance on Juliet Java (with an F1 score over 0.75 across CWEs) than Juliet C/C++ (with two CWEs reporting F1 scores under 0.4).
-On Juliet Java, CodeQL re- ports the highest performance on CWE-78 with an F1 score of 0.92.
-Interestingly, CWE-78 reports the worst results on the other two datasets (F1 score of 0.67 with OWASP and 0.03 with Juliet C/C++).
-[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_47137_highlighted.txt
-In Listing LABEL:fig:julietcpp47137, the model incorrectly reasons about the sanitization by overlooking the & which would cause both the expressions to be evaluated.
-The identified sources and sinks are correct, however.
-| Dataset | CWE | CodeQL | GPT-4 (CWE-DF) | ||||||
-| A | P | R | F1 | A | P | R | F1 | ||
-| OWASP | 22 | 0.52 | 0.50 | 1.00 | 0.67 | 0.48 | 0.48 | 1.00 | 0.64 |
-| OWASP | 78 | 0.56 | 0.53 | 0.77 | 0.63 | 0.48 | 0.48 | 1.00 | 0.65 |
-| OWASP | 89 | 0.57 | 0.54 | 1.00 | 0.70 | 0.52 | 0.52 | 1.00 | 0.68 |
-| OWASP | 79 | 0.79 | 0.71 | 1.00 | 0.83 | 0.65 | 0.59 | 1.00 | 0.74 |
-| Juliet Java | 78 | 0.94 | 0.86 | 1.00 | 0.92 | 0.57 | 0.47 | 1.00 | 0.64 |
-| Juliet Java | 89 | 0.88 | 0.85 | 1.00 | 0.92 | 0.69 | 0.68 | 1.00 | 0.81 |
-| Juliet Java | 476 | 0.76 | 0.76 | 0.76 | 0.76 | 0.79 | 0.75 | 0.88 | 0.81 |
-| Juliet Java | 190 | 0.97 | 1.00 | 0.92 | 0.96 | 0.71 | 0.62 | 0.97 | 0.75 |
-| Juliet C/C++ | 78 | 0.52 | 1.00 | 0.04 | 0.08 | 0.50 | 0.50 | 1.00 | 0.67 |
-| Juliet C/C++ | 476 | 0.80 | 0.95 | 0.75 | 0.84 | 0.71 | 0.75 | 0.86 | 0.80 |
-| Juliet C/C++ | 416 | 0.63 | 1.00 | 0.26 | 0.41 | 0.52 | 0.51 | 0.97 | 0.67 |
-| Juliet C/C++ | 190 | 0.93 | 0.99 | 0.87 | 0.93 | 0.68 | 0.60 | 0.98 | 0.74 |
-A.7. Qualitative analysis of GPT-4 responses
-We first present examples where the dataflow analysis from the CWE-DF prompt is useful.
-Consider the code snippet in Listing 4.
-In this snippet, the variable dir is indirectly being used to create a directory via the dirToCreate variable.
-GPT-4 correctly identifies that this path is not sanitized and could be used to create a directory in otherwise restricted locations.
-This could lead to CWE-22 (path traversal) as is rightly predicted by GPT-4.
-In Listing 5, the method write takes an array b and integers offset and length as inputs.
-The expression offset + length > b.length can lead to an Integer Overflow (CWE-190) if offset + length exceeds the maximum integer value and this is correctly identified by GPT-4.
-Finally, in Listing 6, there are multiple array accesses that could lead to CWE-125 (Out-of-bounds read).
-GPT-4 correctly identifies that the expression ciphertextOffset + length + index could exceed the size of the ciphertext array, leading to this vulnerability.
-A.8. CWE-125 (Out-of-bounds Read in Java vs. C++)
-Listing 6 presents an example of Out-of-bounds read vulnerability in CVEFixes Java which GPT-4 predicts correctly.
-Codellama-34B also correctly predicts that this snippet is vulnerable.
-Listing 7 presents an example from CVEFixes C/C++ where Out-of-bounds read vulnerability is not possible but this is not evident from context.
-Codellama-34B’s explanation seems plausible given that little is known about the pointers in question.